### 17 Jan 2024 v3

# Mean-field underdamped Langevin dynamics and its spacetime discretization

Abstract
We propose a new method called the N-particle underdamped Langevin algorithm for optimizing a special class of non-linear functionals defined over the space of probability measures. Examples of problems with this formulation include training mean-field neural networks, maximum mean discrepancy minimization and kernel Stein discrepancy minimization. Our algorithm is based on a novel spacetime discretization of the mean-field underdamped Langevin dynamics, for which we provide a new, fast mixing guarantee. In addition, we demonstrate that our algorithm converges globally in total variation distance, bridging the theoretical gap between the dynamics and its practical implementation.
Qiang ${}^{\star}$ fuqiang7@mail2.sysu.edu.cn Ashia ${}^{\dagger}$ ashia@mit.edu
${}^{\star}$chool of Mathematics, Sun Yat-sen University
${}^{\dagger}$epartment of Electrical Engineering and Computer Science, MIT
1 Introduction
Many problems in data science can be posed as an entropy-regularized mean-field optimization (EMO) problem. This includes classical applications such as training mean-field neural networks (Mei et al., 2018; Hu et al., 2019; Chizat, 2022; Nitanda et al., 2022; Chen et al., 2022; Claisse et al., 2023; Suzuki et al., 2023), density estimation via maximum mean discrepancy (MMD) minimization (Gretton et al., 2006; Arbel et al., 2019; Chizat, 2022; Suzuki et al., 2023) and sampling via kernel Stein discrepancy (KSD) minimization (Liu et al., 2016; Chwialkowski et al., 2016; Suzuki et al., 2023). A popular approach for solving an EMO problem is the mean-field Langevin dynamics. When the problem can be expressed as optimizing a convex functional, Hu et al. (2019) have demonstrated that mean-field Langevin dynamics converge to a minimizer asymptotically. Additionally, if the problem satisfies a uniform logarithmic Sobolev inequality, several studies have established that this convergence occurs exponentially quickly (Chizat, 2022; Nitanda et al., 2022; Chen et al., 2022). However, implementing the mean-field Langevin dynamics is not a straightforward task; to arrive at a practical algorithm, both spatial and temporal discretizations of the dynamics is required.
Towards this end, a time-discretization of the mean-field Langevin dynamics was analyzed by Nitanda et al. (2022) who extend the interpolation argument introduced by Vempala and Wibisono (2019) to a non-linear Fokker-Planck equation. They establish the non-asymptotic convergence of the discrete-time process in the energy gap. A space-discretization was studied by Chen et al. (2022) who show that the finite-particle approximation to the density of the mean-field Langevin dynamics (referred to as a uniform-in-time propagation of chaos) converges to the minimizer of EMO problem exponentially fast, with a bias related to the number of particles. More practically, Suzuki et al. (2023) analyze a spacetime discretization of the mean-field Langevin dynamics and establish the non-asymptotic convergence of the resulting algorithm to a biased limit related to both the number of particles used and stepsize. Their analysis applies to several important learning problems and improves the results of the standard gradient Langevin dynamics.
A natural candidate method for finding solutions to EMO problems faster is the mean-field underdamped Langevin dynamics. These dynamics resemble several techniques for adding momentum to gradient descent in optimization, many of which are known to result in provably faster convergence in a variety of settings (Nesterov, 1983; Wilson et al., 2016; Laborde and Oberman, 2020; Fu et al., 2023). Moreover, training neural networks using momentum-based gradient descent is considered effective in several applications (Sutskever et al., 2013; Kingma and Ba, 2014; Ruder, 2016). Kazeykina et al. (2020) and Chen et al. (2023) confirm that the naive spacetime discretization of the mean-field underdamped Langevin dynamics have impressive empirical performance when compared to naive discretization of the mean-field Langevin dynamics on applications such as training mean-field neural networks.
In an effort to better understand the behavior of the mean-field underdamped Langevin dynamics, Chen et al. (2023) establish its exponential rate of convergence and propose a finite-particle system that achieves a similar rate under assumptions that are easy to verify for the problem of training mean-field neural networks. In addition, Chen et al. (2023) implement an Euler-Maruyama discretization of the finite-particle system and show that it performs empirically faster when compared with the spacetime discretization of the mean-field Langevin dynamics in training a toy neural network model. However, spacetime discretization of the mean-field underdamped Langevin dynamics are not yet theoretically well understood. Furthermore, the rate obtained by Chen et al. (2023) for the dynamics does not resemble an accelerated rate when compared with Nitanda et al. (2022) and Chen et al. (2022) who analyze the overdamped setting.
A summary of our work
A remaining question is whether we can theoretically characterize the behavior of an implementable algorithm based on discretizing the mean-field underdamped dynamics. If there is a limiting bias, how does it scale with the number of particles and other problem parameters? Ideally, this characterization would give a sharper rate of convergence than Suzuki et al. (2023)’s spacetime discretization of the mean-field Langevin dynamics, suggesting there might be an advantage to adding momentum in the mean-field setting (at least in the worst case). In this paper, we introduce a fast implementable algorithm for solving EMO problems based on the mean-field underdamped Langevin dynamics. We prove that our proposed algorithm converges to a small limiting bias under a set of assumptions that subsumes many problems of interest. In particular, our contributions are summarized as follows.
  1. 1.
    We sharpen the convergence guarantees for both the mean-field underdamped Langevin dynamics and its space-discretization under the same set of assumptions utilized by Chen et al. (2023) (formalized in Theorems 3.1 and 3.2 and summarized in Table 1).
  2. 2.
    We introduce a novel spacetime discretization of the mean-field underdamped Langevin dynamics which we call the N-particle underdamped Langevin algorithm (Algorithm 1). We show the global convergence of our discretization in terms of the total variation (TV) distance. Importantly, our results improve on Suzuki et al. (2023)’s analysis of the spacetime discretization of the mean-field Langevin dynamics, and the additional assumptions needed to prove our result are satisfied by the same examples introduced by Suzuki et al. (2023); these results provide theoretical characterizations of momentum methods in several real-world applications including training neural networks, density estimation and sampling.
Organization
The remainder of this work is organized as follows. Section 2 presents the formal definitions and assumptions as well as important related work. Section 3 proposes our main methods and theoretical results. Section 4 discusses the application of our methods to some classical problems. Section 5 describes our numerical experiments verifying the effectiveness of our proposed methods.
2 Preliminaries
We begin by introducing some general notation that will be used throughout this work.
2.1 Notation
[…]
2.2 Background
We consider entropy-regularized mean-field optimization (EMO) problems, described by
$\min_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}F(\mu)+\lambda\text{Ent}(\mu),$
where $\lambda>0$ and $F$ is a non-linear functional and $\text{Ent}(\mu)=\int\mu\log\mu$ is the negative entropy of $\mu$. The gradient flow structure in 2-Wasserstein metric of the EMO objective is called the mean-field Langevin dynamics,
$\mathrm{d}x_{t}=-D_{\mu}F(\mu_{t},x_{t})\mathrm{d}t+\sqrt{2\lambda}\mathrm{dB}_{t}.$ (MLD)
Hu et al. (2019) show that MLD converges to the solution of EMO problem, $\mu_{*}(x)\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}(\mu_{*},x)\right),$ under mild conditions. We study the mean-field underdamped Langevin dynamics,
$\mathrm{d}x_{t}$ $=v_{t}\mathrm{d}t$ (MULD)
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-D_{\mu}F(\mu_{t}^{x},x_{t})\mathrm{d}t+\sqrt{2\lambda\gamma}\mathrm{dB}_{t},$
and show a new sharp mixing-time bound; here, $\gamma>0$ is the damping coefficient; $\mu_{t}(x,v)=\text{Law}(x_{t},v_{t})$ and $\mu_{t}^{x}\triangleq\text{Law}(x_{t})=\int\mu_{t}(x,v)\mathrm{d}v$ is the $x$-marginal of $\mu_{t}$. The limiting distribution of MULD is the solution to the augmented EMO problem,
$\min_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{2d})}F(\mu^{x})+\lambda\text{Ent}(\mu)+\int\frac{1}{2}\|v\|^{2}\mu(\mathrm{d}x\mathrm{d}v),$ (1)
where a momentum term is added to the EMO objective. To obtain the solution of the EMO problem, it suffices to marginalize the minimizer of (1) given by, $\mu_{*}(x,v)\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}(\mu^{x}_{*},x)-\frac{1}{2\lambda}\|v\|^{2}\right).$ This work also sharpens the analysis of the space-discretization of MULD introduced by Chen et al. (2023), which we refer to as the $N$-particle underdamped Langevin dynamics,
$\mathrm{d}x_{t}^{i}$ $=v_{t}^{i}\mathrm{d}t$ (NULD)
$\mathrm{d}v_{t}^{i}$ $=-\gamma v_{t}^{i}\mathrm{d}t\vspace{-10cm}-D_{\mu}F(\mu_{\textbf{x}_{t}},x^{i}_{t})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}^{i}_{t};$
$i=1,...,N$ where $\mu_{\textbf{x}}\triangleq\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}$; $\mu^{i}_{t}\triangleq\text{Law}(x_{t}^{i},v_{t}^{i})$ and $(\mathrm{B}_{t}^{i})_{i=1}^{N}$ are $d$-dimensional Brownian motions independent of $i$. Note, a time-discretization of NULD is necessary to run the method on a machine. We explore two time-discretization techniques in this paper.
To motivate our algorithm as a time-discretization of NULD, we review discretizations of the underdamped Langevin dynamics (ULD), which is a special case of MULD where $F(\mu)=\int V(x)\mu(\mathrm{d}x)$ is a linear functional of $\mu$:
$\mathrm{d}x_{t}$ $=v_{t}\mathrm{d}t$ (ULD)
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-\nabla V(x_{t})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}_{t}.$
The ULD was first studied in Kolmogoroff (1934) and Hörmander (1967). Under functional inequalities such as Poincaré’s inequality, the convergence guarantee of the ULD was first shown by Villani (2001) and Villani (2009) using a hypocoercivity approach, but without capturing the acceleration phenomenon compared to the overdamped Langevin dynamics. Cao et al. (2023) are the first to achieve the acceleration of the ULD with convex $V$ in $\chi^{2}$-divergence. They prove that when $\mathscr{C}_{\textsf{LSI}}\ll 1$, the decaying rate of ULD is $O(\sqrt{\mathscr{C}_{\textsf{LSI}}})$ whereas the decaying rate of the overdamped Langevin dynamics is $O(\mathscr{C}_{\textsf{LSI}})$.
A discretization of ULD is referred to as an underdamped Langevin Monte Carlo (ULMC) algorithm. There are various discretization schemes proposed for implementing ULD. The Euler-Maruyama (EM) discretization of ULD,
$x_{k+1}$ $=x_{k}+hv_{k},$ (EM-ULMC)
$v_{k+1}$ $=(1-\gamma h)v_{k}-h\nabla V(x_{k})+\sqrt{2\gamma h}\xi_{k},$
for stepsize $h$, $\xi_{k}\sim\mathcal{N}(0,I_{d})$ and $t\in[kh,(k+1)h]$, has been well-studied and shown to incur the largest error (Kloeden et al., 1995; Platen and Bruti-Liberati, 2010). Recently, however, several works derive the convergence of a new ULMC algorithm in Wasserstein distance (Cheng et al., 2018), KL divergence (Ma et al., 2021) and Rényi divergence (Zhang et al., 2023) using a more precise discretization scheme, often referred to as the left point method (LPM) (Hu et al., 2021) and given by
$\mathrm{d}x_{t}$ $=v_{t}\mathrm{d}t,$ (LPM-ULMC)
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-\nabla V(x_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}_{t},$
for $t\in[kh,(k+1)h]$, which is a group of linear stochastic differential equations (SDE) that can be exactly integrated. Unlike the EM scheme, LPM only fixes the drift term in each small interval. Leimkuhler et al. (2023) show that the LPM incurs weaker stepsize restriction when compared with EM. Other discretization schemes are proposed in Shen and Lee (2019); Li et al. (2019); He et al. (2020); Foster et al. (2021); Monmarché (2021); Foster et al. (2022); Johnston et al. (2023), whose convergence guarantee are obtained in Wasserstein distance without achieving better dependence on terms such as the smoothness and LSI constants. In this work, we show that LPM can be applied to discretize both MULD and NULD to produce an implementable algorithm that converges quickly.
2.3 Definitions and assumptions
For each method considered, we study their behavior in settings where the objective satisfies a Log-Sobolev Inequality (LSI).
Definition 1 (LSI).
A measure $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$ satisfies Log-Sobolev Inequality (LSI) with parameter $\mathscr{C}_{\textsf{\em LSI}}>0$, if for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$
[…]
We introduce three essential assumptions on $F$ given by Chen et al. (2023) for establishing the non-asymptotic convergence of the MULD and NULD.
Assumption 2.1 (Convexity).
The functional $F$ is convex, which satisfies
$F(t\mu_{1}+(1-t)\mu_{2})\leq tF(\mu_{1})+(1-t)F(\mu_{2})$ (3)
for any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $t\in[0,1]$.
Assumption 2.2 ($\mathscr{L}$-smoothness).
The functional $F$ is $\mathscr{L}$-smooth; that is, the intrinsic derivative exists and satisfies
$\|D_{\mu}F(\mu_{1},x_{1})-D_{\mu}F(\mu_{2},x_{2})\|\leq\mathscr{L}(W_{1}(\mu_{1},\mu_{2})+\|x_{1}-x_{2}\|)$ (4)
for any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x_{1},x_{2}\in\mathbb{R}^{d}$. We assume $1\leq\mathscr{L}<\infty$ throughout this paper.
Assumption 2.3 (LSI).
The proximal Gibbs distribution
$\hat{\mu}(x,v)=\exp\left(-\frac{\delta F}{\delta\mu}(\mu^{x},x)-\frac{1}{2}\|v\|^{2}\right)/Z(\mu)$ (5)
satisfies LSI with constant $0<\mathscr{C}_{\textsf{\em LSI}}\leq 1$ uniformly in $\mu$.
Note that if $\hat{\mu}^{x}(x)\propto\exp(-\frac{\delta F}{\delta\mu}(\mu^{x},x))$ satisfies LSI with constant $\rho>0$, Assumption 2.3 is satisfied with the choice $\mathscr{C}_{\textsf{LSI}}=\frac{1}{2}\land\rho$. And we refer our readers to Chen et al. (2022, 2023); Suzuki et al. (2023) for the verification of Assumptions 2.1 and 2.3 in a variety of settings. Suzuki et al. (2023) consider a weaker smoothness assumption than Assumption 2.2 where they use $W_{2}$ distance. They verify $W_{2}$ smoothness for three examples including training mean-field neural networks, MMD minimization and KSD minimization, whereas Chen et al. (2022) verify $W_{1}$ smoothness only for training mean-field neural networks. In this paper, we verify $W_{1}$ smoothness in Assumption 2.2 for the other two examples. Beyond Assumptions 2.1-2.3, we introduce four additional assumptions that are sufficient for our spacetime discretization analysis.
Assumption 2.4 (Bounded Gradient).
The intrinsic derivative of $F$ satisfies (where $\mathscr{L}>0$)
$\|D_{\mu}F(\mu,x)\|\leq\mathscr{L}(1+\|x\|).$ (6)
Notably, Suzuki et al. (2023) assume that $F$ can be decomposed as $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ where $\|D_{\mu}U(\mu,x)\|\leq R$ for any $\mu\in\mathcal{P}(\mathbb{R}^{d})$, $x\in\mathbb{R}^{d}$, and where $r(x)$ is $\lambda_{2}$-smooth with $\nabla r(0)=0$ in order to establish the convergence of their spacetime discretization of MLD. Thus, their assumption, $\|D_{\mu}F(\mu,x)\|\leq\|D_{\mu}U(\mu,x)\|+\|\nabla r(x)\|\leq R+\lambda_{2}\|x\|$, implies Assumption 2.4 with the choice $\mathscr{L}\geq\max\{R,\lambda_{2}\}$. The next three assumptions are needed for bounding the second moment of the iterates $(x_{t},v_{t})$ and $(x_{t}^{i},v_{t}^{i})$ along MULD and NULD, which is crucial for the establishment of our discrete-time convergence:
Assumption 2.5.
For all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the proximal Gibbs distribution (5) satisfies $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=O(d).$
Assumption 2.6.
The functional $F$ and the initialization measure $\mu_{0}$ satisfy $F(\mu^{x}_{0})=O(\mathscr{L}d).$
Assumption 2.7.
The functional $F$ and the initialization $\mu_{0}^{N}$ satisfy $\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}F(\mu_{\textbf{x}_{0}})=O(\mathscr{L}d).$
While Assumptions 2.5-2.7 are sufficient, they may not be necessary for the iterates to be bounded. Nevertheless, we argue these assumptions are not too restrictive by providing several examples satisfying them in Section 4.
2.4 Related work
Techniques for establishing the continuous-time convergence of the mean-field underdamped systems and their space-discretization (N-particle systems) are centered around coupling and hypocoercivity, also known as functional approaches (Villani, 2009). The coupling approach generally constructs a joint probability of the mean-field and N-particle systems to make the analytic comparison between them. Based on coupling approaches, Guillin et al. (2022); Bolley et al. (2010); Bou-Rabee and Schuh (2023) obtain the continuous-time convergence of the underdamped systems with mean-field interaction and its space-discretization. Duong and Tugaut (2018); Kazeykina et al. (2020) study the ergodicity of the MULD without a quantitative rate. Under the setting of small mean-field dependence, Kazeykina et al. (2020) obtain the exponential contraction using coupling techniques in Eberle et al. (2019a, b). However, the small mean-field dependence is not satisfied in many applications. The functional approach (hypocoercivity) generally constructs appropriate Lyapunov functionals and studies how their values change along the dynamics. Based on hypocoercivity, Monmarché (2017); Guillin et al. (2021); Guillin and Monmarché (2021); Bayraktar et al. (2022) establish the exponential convergence of the mean-field underdamped systems and its propagation of chaos by constructing a suitable Lyapunov functional. Nevertheless, most of the works above only consider specific settings of MULD such as singular interactions and two-body interactions, which restricts the application to real-world problems. Setting $\gamma=1$, Chen et al. (2023) establish the exponential convergence of MULD and NULD using the hypocoercivity technique in Villani (2009). Under Assumptions 2.1-2.3, they derive the convergence without restricting the size of interactions, which subsumes many settings above. Notably, the techniques of our Theorems 3.1 and 3.2 are adopted from Chen et al. (2023) based on hypocoercivity where we consider other choices of $\gamma$ to improve the decaying rate of MULD and NULD established in Chen et al. (2023).
3 N-particle underdamped Langevin algorithm
Our first step is to establish the global convergence of the mean-field underdamped Langevin algorithm,
$\mathrm{d}x_{t}$ $=v_{t}\mathrm{d}t,$ (MULA)
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-D_{\mu}F(\mu^{x}_{k},x_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}_{t},$
for stepsize $h$, $t\in[kh,(k+1)h]$ and $k=1,...,K$. Note that MULA is a time-discretization of the MULD, where each step requires integrating from $t=kh$ to $t=(k+1)h$ for stepsize $h$. MULA is intractable to implement in most instances given we do not often have access to $\mu_{k}^{x}$ per iteration. This prompts us to consider the particle approximation which uses $\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{k}^{i}}$ to approximate $\mu_{k}^{x}$:
$\mathrm{d}x_{t}^{i}$ $=v^{i}_{t}\mathrm{d}t,$ (7)
$\mathrm{d}v_{t}^{i}$ $=\gamma v^{i}_{t}\mathrm{d}t-D_{\mu}F(\mu_{\textbf{x}_{k}},x^{i}_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}^{i}_{t}$
for stepsize $h$, $t\in[kh,(k+1)h]$, $i=1,...,N$, $k=1,\dots$ and $\mu_{\textbf{x}_{k}}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x^{i}_{k}}$. Integrating the particle system  (7) from $t=kh$ to $t=(k+1)h$ for stepsize $h$ and $i=1,...,N$, we obtain our proposed Algorithm 1 which we refer to as the N-particle underdamped Langevin algorithm.
Algorithm 1 N-particle underdamped Langevin algorithm (NULA)
0:  $F$ satisfies Assumptions 2.1-2.5 and 2.7
1:  Initialize $\textbf{x}_{0}=(x_{0}^{1},...,x_{0}^{N}),\,\textbf{v}_{0}=(v_{0}^{1},...,v_{0}^{N})$, $h,\gamma$ Specify $\varphi_{2}=\exp(-\gamma h),\,\varphi_{0}=(1/\gamma)(1-\varphi_{2}),\varphi_{1}=(1/\gamma)(h-\varphi_{0}),\Sigma^{11}=(2/\gamma)(h-2\varphi_{0}+(1/2\gamma)(1-\varphi_{2}^{2})),\Sigma^{12}=(1/\gamma)(1-\varphi_{2})^{2},\,\Sigma^{22}=1-\varphi_{2}^{2}$
2:  for $k=0,...,K-1$ do
3:     for $i=1,...,N$ do
4:        $\left[\begin{array}[]{c}(\mathrm{B}_{k}^{i})^{x}\\ (\mathrm{B}_{k}^{i})^{v}\end{array}\right]\sim\mathcal{N}\left(0,\left[\begin{array}[]{cc}\Sigma^{11}I_{d}&\Sigma^{12}I_{d}\\ \Sigma^{12}I_{d}&\Sigma^{22}I_{d}\end{array}\right]\right)$
5:        $x^{i}_{k+1}=x^{i}_{k}+\varphi_{0}\,v^{i}_{k}-\varphi_{1}\,D_{\mu}F(\mu_{\textbf{x}_{k}},x^{i}_{k})+(\mathrm{B}_{k}^{i})^{x}$
6:        $v^{i}_{k+1}=\varphi_{2}\,v^{i}_{k}-\varphi_{0}\,D_{\mu}F(\mu_{\textbf{x}_{k}},x^{i}_{k})+(\mathrm{B}_{k}^{i})^{v}$
7:     end for
8:  end for
9:  return  $(x^{1}_{K},...,x^{N}_{K})$
The update parameters of Algorithm 1, $\varphi_{0},\,\varphi_{1},\,\varphi_{2}$ and $(\mathrm{B}_{k}^{i})^{x},\,(\mathrm{B}_{k}^{i})^{v}$, are functions of $\gamma$ and stepsize $h$. Thus we need to specify the value of $\gamma$ and $h$ to compute the update parameters and initialize $(\textbf{x}_{0},\,\textbf{v}_{0})$ before running the algorithm.
3.1 Convergence analysis
We begin by leveraging Theorem 2.1 and Theorem 2.3 from Chen et al. (2023) analyzing the continuous-time dynamics using entropic hypocoercivity. Throughout, let
$S=\left(\begin{array}[]{cc}{1}/{\mathscr{L}}&{1}/{\sqrt{\mathscr{L}}}\\ {1}/{\sqrt{\mathscr{L}}}&2\end{array}\right)\otimes I_{d}.$
Both proofs utilize a Lyapunov functional similar to the one introduced by Chen et al. (2023). In particular, Theorem 3.1 is established by showing the following functional,
$\mathcal{E}(\mu)$ $\triangleq\mathcal{F}(\mu)+\textsf{FI}_{S}(\mu\|\hat{\mu}),\,\text{where}$ (8)
$\mathcal{F}(\mu)$ $\triangleq F(\mu^{x})+\int\frac{1}{2}\|v\|^{2}\mu(\mathrm{d}x\mathrm{d}v)+\text{Ent}(\mu),$
is decaying along the trajectory of the dynamics. Our second Theorem 3.2 establishes convergence of the N-particle system NULD. Using the notation $\textbf{x}=(x^{1},...,x^{N})$, $\textbf{v}=(v^{1},...,v^{N})$ and $\mu^{N}=\otimes_{i=1}^{N}\mu^{i}$, we take $\mu_{*}^{N}$ to be the limiting distribution of NULD satisfying $\mu_{*}^{N}\propto\exp\left(-NF(\mu_{\textbf{x}})-\frac{1}{2}\|\textbf{v}\|^{2}\right)$. With these definitions, we obtain our guarantee by showing the functional,
$\mathcal{E}^{N}(\mu^{N})$ $\triangleq\mathcal{F}^{N}(\mu^{N})+\textsf{FI}^{N}_{S}(\mu^{N}\|\mu_{*}^{N}),\,\text{where}$ (9)
$\textsf{FI}^{N}_{S}(\mu^{N}\|\mu_{*}^{N})$ $\triangleq\sum_{i=1}^{N}\mathbb{E}_{\mu^{N}}\Big{\|}S^{1/2}\nabla_{i}\log\frac{\mu^{N}}{\mu_{*}^{N}}\Big{\|}^{2},\,\text{and}$
$\mathcal{F}^{N}(\mu)$ $\triangleq\int NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2}\mu^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu^{N}),$
is decaying along the trajectory of NULD, where $\nabla_{i}=(\nabla_{x^{i}},\nabla_{v^{i}})^{\mathsf{T}}$.
Theorem 3.1 (Mean-field underdamped Langevin dynamics).
If Assumptions 2.1-2.3 hold, $\mu_{0}$ has finite second moment, finite entropy and finite Fisher information, then the law $\mu_{t}$ of the MULD with $\gamma=\sqrt{\mathscr{L}}$ and $\mathcal{E}$ defined in (8) satisfy,
$\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*})\leq(\mathcal{E}(\mu_{0})-\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{\em LSI}}}{3\sqrt{\mathscr{L}}}t\right).$
Theorem 3.2 (N-particle underdamped Langevin dynamics).
If Assumptions 2.1-2.3 hold, $\mu_{0}^{N}$ has finite second moment, finite entropy, finite Fisher information, and $N\geq(\mathscr{L}/\mathscr{C}_{\textsf{\em LSI}})\left(32+24\mathscr{L}/\mathscr{C}_{\textsf{\em LSI}}\right)$, then the joint law $\mu_{t}^{N}$ of the NULD with $\gamma=\sqrt{\mathscr{L}}$ and $\mathcal{E}^{N}$ defined in (9) satisfy
$\frac{1}{N}\mathcal{F}^{N}(\mu_{t}^{N})-\mathcal{F}(\mu_{*})\leq\frac{\mathcal{E}_{0}^{N}}{N}\exp\left(-\frac{\mathscr{C}_{\textsf{\em LSI}}}{6\sqrt{\mathscr{L}}}t\right)+\frac{\mathcal{B}}{N},$
where $\mathcal{B}=\frac{60\mathscr{L}d}{\mathscr{C}_{\textsf{\em LSI}}}+\frac{36\mathscr{L}^{2}d}{\mathscr{C}^{2}_{\textsf{\em LSI}}}$, $\mathcal{E}_{0}^{N}\triangleq\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*})$.
Theorem 3.2 implies the non-uniform-in-N convergence of NULD, which incorporates a bias term involving $N$ due to the particle approximation. Our proof technique mirrors that of Chen et al. (2023); however, we demonstrate that a faster convergence and smaller bias in Theorem 3.2 can be achieved by choosing $\gamma=\sqrt{\mathscr{L}}$ instead of $\gamma=1$ (see Table 1).
Our main theorems analyzes the convergence of the discrete-time processes and their mixing time guarantees to generate an $\epsilon$-approximate solution in TV distance with the specific choice of initialization, damping coefficient $\gamma$, and stepsize $h$.
Theorem 3.3 (Mean-field underdamped Langevin algorithm).
In addition to the assumptions specified in Theorems 3.1, let Assumptions 2.4-2.6 hold. Denote $\bar{\mu}_{K}$ the law of $(x_{K},v_{K})$ of the MULA and $\kappa\triangleq\mathscr{L}/\mathscr{C}_{\sf{LSI}}$. Then in order to ensure $\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{\em TV}}\leq\epsilon$, it suffices to choose $\gamma=\sqrt{\mathscr{L}}$, $\bar{\mu}_{0}=\mathcal{N}(0,I_{2d})$, and
$h=\widetilde{\Theta}\left(\frac{\mathscr{C}_{\textsf{\em LSI}}\epsilon}{\mathscr{L}^{3/2}d^{1/2}}\right),\ K=\widetilde{\Theta}\left(\frac{\kappa^{2}d^{1/2}}{\epsilon}\right).$
A similar guarantee can be stated for the $N$-particle system (7) with the additional requirement that the number of particles scale according to the dimension of the problem and problem parameters.
Theorem 3.4 (N-particle underdamped Langevin algorithm).
In addition to the assumptions specified in Theorem 3.2, let Assumptions 2.4, 2.5 and 2.7 hold. Denote $\bar{\mu}^{i}_{K}$ the law of $(x_{K}^{i},v_{K}^{i})$ of the NULA for $i=1,...,N$ and $\kappa\triangleq\mathscr{L}/\mathscr{C}_{\mathsf{LSI}}$. Then in order to ensure $\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}_{K}^{i}-\mu_{*}\|_{\textsf{\em TV}}\leq\epsilon$, it suffices to choose $\gamma=\sqrt{\mathscr{L}}$, $\bar{\mu}^{N}_{0}=\mathcal{N}(0,I_{2Nd})$,
$h=\widetilde{\Theta}\left(\frac{\mathscr{C}_{\textsf{\em LSI}}\epsilon}{\mathscr{L}^{3/2}d^{1/2}}\right),\ K=\widetilde{\Theta}\left(\frac{\kappa^{2}d^{1/2}}{\epsilon}\right),$
and the number of particles $N=\Theta\left(\kappa^{2}d/\epsilon^{2}\right).$
Discretization Method # of particles Mixing time
Time-discretizations MLA (Nitanda et al., 2022) * $\widetilde{\Theta}\Big{(}\kappa^{2}\mathscr{L}d/\epsilon^{2}\Big{)}$
LPM-ULMC (Zhang et al., 2023) * $\widetilde{\Theta}\Big{(}\kappa^{3/2}d^{1/2}/\epsilon\Big{)}$
MULA (Ours) * $\widetilde{\Theta}\left(\kappa^{2}d^{1/2}/\epsilon\right)$
Space-discretizations NULD (Chen et al. (2023)) ${\Theta}\left(\kappa^{2}\mathscr{L}d/\epsilon^{2}\right)$ $\widetilde{\Theta}\Big{(}\kappa\Big{)}$
NULD (Ours) $\Theta\Big{(}\kappa^{2}d/\epsilon^{2}\Big{)}$ $\widetilde{\Theta}\Big{(}\kappa/\mathscr{L}^{1/2}\Big{)}$
spacetime discretizations NLA (Suzuki et al., 2023) ${\Theta}\left(\kappa\mathscr{L}^{3}d/\epsilon^{2}\right)$ $\widetilde{\Theta}\left(\kappa^{2}\mathscr{L}d/\epsilon^{2}\right)$
NULA (Ours) $\Theta\left(\kappa^{2}d/\epsilon^{2}\right)$ $\widetilde{\Theta}\Big{(}\kappa^{2}d^{1/2}/\epsilon\Big{)}$
Table 1: Comparison of algorithms in terms of the mixing time and number of particles to achieve $\epsilon$-approximate solutions in TV distance. $\kappa\triangleq\mathscr{L}/\mathscr{C}_{\textsf{LSI}}$. * represents that we do not need particle approximation for this method.
3.2 Proof sketches
For the continuous-time results, we outline the proof of Theorem 3.1 (and analogously Theorem 3.2) in this section to provide intuition for how choosing $\gamma=\sqrt{\mathscr{L}}$ can improve the decaying rate of MULD. We begin with a review of some notations of hypocoercivity in Villani (2009); Chen et al. (2023):
$A_{t}=\nabla_{v},\ C_{t}=\nabla_{x},\ Y_{t}=(\|A_{t}u_{t}\|,\|A_{t}^{2}u_{t}\|,\|C_{t}u_{t}\|,\|C_{t}A_{t}u_{t}\|)^{\mathsf{T}},$
where $u_{t}=\log\frac{\mu_{t}}{\hat{\mu}_{t}}.$ Inheriting the analysis of Theorem 2.1 in Chen et al. (2023) and Lemma 32 in Villani (2009), we show that for a general $\gamma$, the Lyapunov functional (8) with $S=[s_{ij}]\otimes I_{d}\in\mathbb{R}^{2d\times 2d}$ is decreasing along MULD satisfying
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-Y_{t}^{\mathsf{T}}\mathcal{K}Y_{t},$ (10)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-\lambda_{\textsf{min}}(\mathcal{K})Y_{t}^{\mathsf{T}}Y_{t}$ $\leq-\mathscr{C}_{\textsf{LSI}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{2\lambda_{\textsf{max}}(S)}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu_{t}})$
$\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{6\mathscr{L}}(\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*}))$
Applying Grönwall’s inequality leads to the decaying rate $O(\mathscr{C}_{\textsf{LSI}}/\mathscr{L})$ of MULD ($\gamma=1$) in Chen et al. (2023). If we specify $\gamma=\sqrt{\mathscr{L}}$, we can choose $b=1/\sqrt{\mathscr{L}},\,a=2,\,c=1/\mathscr{L}$ satisfying the criteria. Then we obtain $\lambda_{\textsf{min}}(\mathcal{K})=2/\sqrt{\mathscr{L}}$ and
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-\lambda_{\textsf{min}}(\mathcal{K})Y_{t}^{\mathsf{T}}Y_{t}$ $\leq-\frac{2\mathscr{C}_{\textsf{LSI}}}{\sqrt{\mathscr{L}}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{\lambda_{\textsf{max}}(S)\sqrt{\mathscr{L}}}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu}_{t})$
$\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{3\sqrt{\mathscr{L}}}(\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*}))$
Applying Grönwall’s inequality leads to the improved decaying rate $O(\mathscr{C}_{\textsf{LSI}}/\sqrt{\mathscr{L}})$ of MULD ($\gamma=\sqrt{\mathscr{L}}$) in our Theorem 3.1. We defer the whole proof to Appendix C.
For discretization errors, we outline the proof of Theorem 3.3 (and analogously Theorem 3.4) in this section. Let $(\mu_{t})_{t\geq 0}$ and $(\bar{\mu}_{t/h})_{t\geq 0}$ represent the law of MULD and MULA initialized at $\mu_{0}$. Let $\textbf{Q}_{kh}$ and $\textbf{P}_{kh}$ denote probability measures of MULD and MULA on the space of paths $C([0,kh],\mathbb{R}^{2d})$. Invoking Girsanov’s theorem (Girsanov, 1960; Kutoyants, 2004; Le Gall, 2016) and Assumption 2.2, we can upper bound the pathwise divergence between MULD and MULA in KL divergence for stepsize $h$ and $k=1,...,K$ under Assumptions 2.2 and 2.4:
$\textsf{KL}(\textbf{Q}_{Kh}\|\textbf{P}_{Kh})\lesssim\frac{\mathscr{L}^{4}h^{5}}{\gamma}\sum_{k=0}^{K-1}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{kh}\|^{2}+\frac{\mathscr{L}^{2}h^{3}}{\gamma}\sum_{k=0}^{K-1}\mathbb{E}_{\mathbf{Q}_{Kh}}\|v_{kh}\|^{2}+\frac{\mathscr{L}^{4}h^{5}K}{\gamma}+{\mathscr{L}^{2}h^{4}Kd}$ (11)
The derivation of (11) is similar to that of Zhang et al. (2023); they establish the discretization error of LPM-ULMC in $q$-th order Rényi divergence ($q\in[1,2)$), which has KL divergence as a special case ($q=1$). Their smoothness assumption on the potential function $V$ is $(\mathscr{L},s)$-weak smoothness, which recovers $\mathscr{L}$-smoothness when $s=1$. We use many similar techniques of bounding the discretization error to those of Zhang et al. (2023). Their Lemma 26 can be generalized to our Lemma 7 in the mean-field setting, which describes an intermediate process of deriving (11). Applying the data processing inequality, we can upper bound the KL divergence between the time marginal laws of the iterates by KL divergence between path measures:
$\textsf{KL}(\mu_{T}\|\bar{\mu}_{K})\leq\textsf{KL}(\textbf{Q}_{Kh}\|\textbf{P}_{Kh}),$
where $T=Kh$. To uniformly upper bound the right-hand side of (11), we need to uniformly upper bound $\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{kh}\|^{2}$ and $\mathbb{E}_{\mathbf{Q}_{Kh}}\|v_{kh}\|^{2}$. The techniques of uniformly upper bounding the iterates of ULD in Zhang et al. (2023) rely on $\chi^{2}$-convergence guarantee of ULD. However, $\chi^{2}$-convergence is not established for MULD by previous works. We thus utilize different techniques to uniformly upper bound the iterates of MULD and NULD. More specifically, we have
$\mathbb{E}_{\textbf{Q}_{T}}\|(x_{t},v_{t})\|^{2}=W_{2}^{2}(\mu_{t},\delta_{0})\lesssim\underbrace{W_{2}^{2}(\mu_{t},\mu_{*})}_{\textsf{I}}+\underbrace{W_{2}^{2}(\mu_{*},\delta_{0})}_{\textsf{II}},\ t\in[0,T],$
where $\delta_{0}$ is Dirac measure on $\textsf{0}\in\mathbb{R}^{2d}$, and II is the second moment of $\mu_{*}$ denoted by $\textbf{m}_{2}^{2}$. Now we need to upper bound I. Under Assumption 2.3, $\mu_{*}$ satisfies LSI implying Talagrand’s inequality: $\textsf{I}\lesssim\textsf{KL}(\mu_{t}\|\mu_{*})/\mathscr{C}_{\textsf{LSI}}$. Under Assumptions 2.1 and 2.2, Lemma 4.2 in Chen et al. (2023) establishes the following relation between KL divergence and energy gap:
$\textsf{KL}(\mu_{t}\|\mu_{*})\leq\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}).$ (12)
Moreover, Kazeykina et al. (2020); Chen et al. (2023) demonstrate that $\mathcal{F}(\mu_{t})$ is decreasing along MULD. According to two conclusions above, I can be bounded as
$\textsf{I}\lesssim\frac{\textsf{KL}(\mu_{t}\|\mu_{*})}{\mathscr{C}_{\textsf{LSI}}}\leq\frac{\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*})}{\mathscr{C}_{\textsf{LSI}}}\leq\frac{\mathcal{F}(\mu_{0})-\mathcal{F}(\mu_{*})}{\mathscr{C}_{\textsf{LSI}}}\leq\frac{\mathcal{F}(\mu_{0})}{\mathscr{C}_{\textsf{LSI}}},$
where the last inequality follows from the assumption that $\mathcal{F}(\mu_{*})\geq 0$. Therefore, under Assumptions 2.5 on $\textbf{m}_{2}^{2}$ and 2.6 on $F(\mu_{0})$, our Lemma 8 establishes the upper bound of $\mathbb{E}_{\textbf{Q}_{T}}\|(x_{t},v_{t})\|^{2}$ in terms of $\mathscr{L},\,\mathscr{C}_{\textsf{LSI}}$ and $d$, which implies the uniform upper bound of $\textsf{KL}(\mu_{T}\|\bar{\mu}_{K})$. Applying Pinsker’s inequality
$\|\bar{\mu}_{K}-\mu_{T}\|_{\textsf{TV}}\lesssim\sqrt{\textsf{KL}(\mu_{T}\|\bar{\mu}_{K})},$
we can convert the discretization error bound in KL divergence to that in TV distance. Combining Pinsker’s inequality and relation (12), we derive the continuous-time convergence of MULD in Theorem 3.1 in TV distance:
$\|\mu_{T}-\mu_{*}\|_{\textsf{TV}}\lesssim\sqrt{\textsf{KL}(\mu_{T}\|\mu_{*})}\leq\sqrt{\mathcal{F}(\mu_{T})-\mathcal{F}(\mu_{*})}.$ (13)
Applying the triangle inequality to $\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}$, the TV distance between the law of MULA at $Kh$ and the limiting distribution of MULD, we obtain the global convergence of MULA:
$\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}$ $\leq\underbrace{\|\bar{\mu}_{K}-\mu_{T}\|_{\textsf{TV}}}_{\mathcal{B}}+\underbrace{\|\mu_{T}-\mu_{*}\|_{\textsf{TV}}}_{\mathcal{V}},$
where $\mathcal{V}$ vanishes exponentially fast as $T\rightarrow\infty$ and $\mathcal{B}$ is a vanishing bias as $h\rightarrow 0$. To ensure $\mathcal{V}+\mathcal{B}\leq\epsilon$, it suffices to choose $T=\widetilde{\Theta}(\sqrt{\mathscr{L}}/\mathscr{C}_{\textsf{LSI}})$ and specify $h,\,K$ as in Theorem 3.3. The whole proof is deferred to Appendix D.
3.3 Discussion of mixing time results
We summarize the convergence results of MULA, NULA and several existing methods including LPM-ULMC, the EM-discretization of MLD (referred to as MLA (Nitanda et al., 2022)), and its finite-particle system (referred to as NLA (Suzuki et al., 2023)) in Table 1. For the mixing time to generate an $\epsilon$-approximate solution in TV distance, our proposed MULA and NULA achieve better dependence on $\mathscr{L}$, $d$ and $\epsilon$ than MLA and NLA, and keep the same dependence on $\mathscr{C}_{\textsf{LSI}}$ as MLA and NLA, which justifies that our methods are fast. For the number of particles, we improve the dependence on $\mathscr{L}$ for NULD ($\gamma=\sqrt{\mathscr{L}}$) when compared with NULD ($\gamma=1$) in Chen et al. (2023) and for NULA when compared with NLA. Particularly, our dependence on the smoothness constant in the number of particle guarantee of NULA is $\Theta(\mathscr{L}^{2})$ whereas the counterpart of NLA is $\Theta(\mathscr{L}^{4})$. However, our dependence on the LSI constant in the number of particle guarantee of NULA is $\Theta(\mathscr{C}^{-2}_{\textsf{LSI}})$ whereas the counterpart of NLA is $\Theta(\mathscr{C}^{-1}_{\textsf{LSI}})$.
Note that Nitanda et al. (2022) consider MLA in the neural network setting where they specifically choose $F$ to be the objective (14) and propose assumptions on $l$, $h$ and $r$. Suzuki et al. (2023) consider NLA in a setting where they specify that $F(\mu)=U(\mu)+\mathbb{E}_{\mu}[r(x)]$ and propose assumptions on $U$ and $r$. Consequently, they use different notations of the smoothness constant and establish the convergence rate in energy gap $\mathcal{F}(\bar{\mu}_{K})-\mathcal{F}(\mu_{*})$ instead of the TV distance. To make a fair comparison, we equivalently translate those smoothness constants into $\mathscr{L}$ and convert convergence rates of MLA and NLA to those in TV distance by relation (12) and Pinsker’s inequality (see Appendix F).
4 Applications of Algorithm 1
In Section 1, we briefly mention three examples on the applications of problem (2.2). In this section, we will show how Algorithm 1 can be applied to these examples by verifying Assumptions 2.1-2.7 for these examples. We defer the proof of verification to Appendix B.
4.1 Training mean-field neural networks
Consider a two-layer neural network in the mean-field regime (with infinite neurons), which can be parametrized as $h(\mu;a)\triangleq\mathbb{E}_{\mu}[h(x;a)],$ where $h(x;a)$ represents a single neuron with trainable parameter $x$ and input $a$ (e.g. $h(x;a)=\sigma(x^{\mathsf{T}}a)$ for activation function $\sigma$); $\mu$ is the probability distribution of the parameter $x$. Given dataset $(a_{i},b_{i})_{i=1}^{n}$, loss function $l$ and regularization function $r$, we can choose $F(\mu)$ in (2.2) to be
$F(\mu)=\frac{1}{n}\sum_{i=1}^{n}l(h(\mu;a_{i}),b_{i})+\lambda^{\prime}\mathbb{E}_{x\sim\mu}[r(x)].$ (14)
Since $l$ is convex and $h(\mu;a)$ is a linear functional of $\mu$, which is also convex, their composition remains convex if $h$ satisfies quadratic growth (Proposition 3.2 in Chen et al. (2022)). The learning of two-layer neural networks is thus transformed into a convex optimization problem in the space of measures, for which the convexity of loss function can be utilized to show the global convergence of the gradient-based optimization methods (Nitanda and Suzuki, 2017; Mei et al., 2018; Chizat and Bach, 2018; Rotskoff and Vanden-Eijnden, 2018b; Sirignano and Spiliopoulos, 2020; Suzuki et al., 2023). The entropy regularization term in (2.2) comes from the noisy gradient descent of $F(\mu)$. While the minimizers of $F(\mu)+\lambda\text{Ent}(\mu)$ and $F(\mu)$ are different, the global convergence analysis of $F(\mu)+\lambda\text{Ent}(\mu)$ requires less stringent assumptions (Rotskoff and Vanden-Eijnden, 2018a; Mei et al., 2019; Hu et al., 2019; Nitanda et al., 2022) and we can control the value of $\lambda$ to make the minimizers of $F(\mu)+\lambda\text{Ent}(\mu)$ close to the minimizers of $F(\mu)$.
The objectives (14) satisfy Assumptions 2.1-2.4 for specific common choices $l$, $h$ and $r$ described in several works  (Nitanda et al., 2022; Chen et al., 2022, 2023; Suzuki et al., 2023). In order to make (14) also satisfy Assumptions 2.5-2.7, assumptions on $l$, $h$ and $r$ proposed in Nitanda et al. (2022) suffice. We only need to additionally specify the initialization without incurring stronger assumptions on $l$, $h$ and $r$. For instance, if we choose $r(x)=\|x\|^{2}/2$; there exists $\mathscr{L}>0$ such that the activation function satisfies $|h(x;a)|\leq\sqrt{\mathscr{L}}$ (also proposed in Suzuki et al. (2023)); and the convex loss function $l$ is quadratic or satisfies $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$ (also proposed in Nitanda et al. (2022)), $F$ will meet Assumption 2.5 with $\lambda^{\prime}\lesssim(2\pi)^{\frac{d}{d-2}}\exp(-\frac{8\mathscr{L}}{d-2})$. Finally, if in addition we assume $l$ is $\sqrt{\mathscr{L}}$-Lipschitz and choose $\mu_{0}=\mathcal{N}(0,I_{2d})$ and $\mu_{0}^{N}=\mathcal{N}(0,I_{2Nd})$, Assumptions 2.6 and 2.7 will be satisfied. We defer other examples and the related verification to Appendix B.
4.2 Density estimation via MMD minimization
Inheriting the definition in Gretton et al. (2006), the maximum mean discrepancy between two probability measures $\mu$ and $\pi$ is defined as $\mathcal{M}(\mu\|\pi)=\iint k(x,x)-2k(x,y)+k(y,y)\mathrm{d}\mu(x)\mathrm{d}\nu(y),$ given a positive definite kernel $k$. Similar to Example 2 in Suzuki et al. (2023), we consider the non-parametric density estimation using the Gaussian mixture model, which can be parametrized as $p(\mu;z)\triangleq\mathbb{E}_{x\sim\mu}[p(x;z)],$ where $p(x;z)$ is the Gaussian density function with mean $x$ and a user-specified variance $\sigma^{2}$. Given a set of samples $\{z_{i}\}_{i=1}^{n}$ from the target distribution $p^{*}$, our goal is to fit $p^{*}$ by minimizing the empirical version of $\mathcal{M}(p(\mu;z)\|p^{*})$, defined as
^$\hat{\mathcal{M}}(\mu)=\iiint p(x;z)p(x^{\prime};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}\mathrm{d}(\mu\times\mu)(x,x^{\prime})-2\int\left(\frac{1}{n}\sum_{i=1}^{n}\int p(x;z)k(z,z_{i})\mathrm{d}z\right)\mathrm{d}\mu(x).$
We can choose $F(\mu)$ in (2.2) to be
$F(\mu)=\hat{\mathcal{M}}(\mu)+\lambda^{\prime}\mathbb{E}_{x\sim\mu}[r(x)],$ (15)
where $\lambda^{\prime}>0$ and $r$ is the regularization function. Suzuki et al. (2023) show that objective (15) satisfies Assumptions 2.1, 2.3 and 2.4 by choosing smooth and light-tailed kernel $k$, such as Gaussian radial basis function (RBF) kernel defined as $k(z,z^{\prime})\triangleq\exp(-\|z-z^{\prime}\|^{2}/2\sigma^{\prime 2})$ for $\sigma^{\prime}>0$. We also verify that objective (15) also satisfies our Assumption 2.2 with the same choice of kernel. With Gaussian RBF kernel $k$ ($\sigma^{\prime}=\sigma$), we demonstrate that objective (15) satisfies Assumptions 2.5-2.7 if we specify $\lambda^{\prime}\lesssim 2\pi$, $r(x)=\|x\|^{2}/2$, $\mu_{0}=\mathcal{N}(0,I_{2d})$ and $\mu_{0}^{N}=\mathcal{N}(0,I_{2Nd})$.
4.3 Kernel Stein discrepancy minimization
Kernel Stein discrepancy (KSD) minimization is a method for sampling from a target distribution $\mu_{*}$ if we have the access to the score function $s_{\mu_{*}}(x)=\nabla\log\mu_{*}(x)$ (Chwialkowski et al., 2016; Liu et al., 2016). For example, if $\mu_{*}(x)\propto\exp(-V(x))$, we need the access of $\nabla V(x)$. For a positive definite kernel $k$, the Stein kernel is defined as
$u_{\mu_{*}}(x,x^{\prime})=s_{\mu_{*}}^{\mathsf{T}}(x)k(x,x^{\prime})s_{\mu_{*}}(x^{\prime})+s_{\mu_{*}}^{\mathsf{T}}(x)\nabla_{x^{\prime}}k(x,x^{\prime})+\nabla_{x}^{\mathsf{T}}k(x,x^{\prime})s_{\mu_{*}}(x^{\prime})+\textsf{tr}(\nabla_{x,x^{\prime}}k(x,x^{\prime})).$
The KSD between $\mu$ and $\mu_{*}$ is defined as
$\textsf{KSD}(\mu)=\iint u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime}).$ (16)
We can choose $F(\mu)$ in (2.2) to be
$F(\mu)=\textsf{KSD}(\mu)+\lambda^{\prime}\mathbb{E}_{x\sim\mu}[r(x)],$ (17)
where $\lambda^{\prime}>0$ and $r$ is the regularization function. Suzuki et al. (2023) show that objective (17) satisfies Assumptions 2.1, 2.3 and 2.4 by choosing light-tailed kernel and assume the score function satisfies
$\max\{\|\nabla\log\mu_{*}(x)\|,\|\nabla^{\otimes 2}\log\mu_{*}(x)\|_{\textsf{op}},\|\nabla^{\otimes 3}\log\mu_{*}(x)\|_{\textsf{op}}\}\leq\mathscr{L}(1+\|x\|).$ (18)
More specifically, if $\mu_{*}\propto\exp(-V)$, the potential function $V$ should satisfies
$\max\{\|\nabla V(x)\|,\|\nabla^{\otimes 2}V(x)\|_{\textsf{op}},\|\nabla^{\otimes 3}V(x)\|_{\textsf{op}}\}\leq\mathscr{L}(1+\|x\|),$
which subsumes many distributions. With the same choice of kernel in Suzuki et al. (2023), we verify that objective (17) also satisfies Assumption 2.2 and satisfies our Assumptions 2.5-2.7 with $\lambda^{\prime}\lesssim(2\pi)^{\frac{d}{d-2}}\exp\left(-\frac{4\mathscr{L}}{d-2}\right)$, $r(x)=\|x\|^{2}/2$, $\mu_{0}=\mathcal{N}(0,I_{d})$ and $\mu_{0}^{N}=\mathcal{N}(0,I_{2Nd})$.
5 Numerical experiments
We verify our theoretical findings by providing empirical support in this section. Our experiment [^1] Code for our experiments can be found at https://github.com/QiangFu09/NULA. is to approximate a Gaussian function $f(z)=\exp(-\|z-m\|^{2}/2d)$ for $z\in\mathbb{R}^{d}$ and unknown $m\in\mathbb{R}^{d}$ by a mean-field two-layer neural network with tanh activation. Consider the empirical risk minimization problem (14) with quadratic loss function $l$, $r(x)=\|x\|^{2}/2$, $d=10^{3}$, $\lambda^{\prime}=10^{-4}$ and $n$ randomly generated data samples from $f(z)$ ($n=100$), described by
$F(\mu)=\frac{1}{2n}\sum_{i=1}^{n}(h(\mu;a_{i})-f(a_{i}))^{2}+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}[\|x\|^{2}].$
$F$ satisfy Assumptions 2.1-2.7 with the choice of $l,\,h,\,r$, and thus we apply Algorithm 1 for minimizing the objective above. Note that the number of neurons in the first hidden layer is equivalent to the number of particles in NULA, and we choose $N\in\{256,512,1024,2048\}$.
Figure 1: Evaluation on NULA, NLA and EM-NULA with different number of particles N where x-axis represents the training epochs and y-axis represents the value of $\frac{1}{2n}\sum_{i=1}^{n}(\frac{1}{N}\sum_{s=1}^{N}h(x^{s};a_{i})-f(a_{i}))^{2}$. Our method often enjoys better performance in the high particle-approximation regime which is consistent with our theoretical findings.
The intrinsic derivative of $F(\mu)$ for the $j$-th particle in our method is given by
${D}_{\mu}F(\mu_{\textbf{x}},x^{j})=\frac{1}{n}\sum_{i=1}^{n}(\frac{1}{N}\sum_{s=1}^{N}h(x^{s};a_{i})-f(a_{i}))\nabla h(x^{j};a_{i})+\lambda^{\prime}x^{j}.$
Note that $\frac{1}{N}\sum_{s=1}^{N}h(x_{s};a)$ is in fact a two-layer neural network with $N$ neurons. Instead of fine-tuning $\gamma$ and stepsize $h$ in NULA, we directly fine-tune the value of $\varphi_{0}$, $\varphi_{1}$ and $\varphi_{2}$ in Algorithm 1 by grid search. For simplifying the computation, we approximate $(\mathrm{B}_{k}^{i})^{x}$ and $(\mathrm{B}_{k}^{i})^{v}$ by $\eta\xi^{x}_{k}$ and $\eta\xi^{v}_{k}$ where $\xi^{x}_{k}$ and $\xi^{v}_{k}$ are independent standard Gaussian, and then we fine-tune the scaling scalar $\eta$. We compare our method (NULA) to NLA with stepsize $h_{1}$ and scaling scalar $\lambda_{1}$ given by,
$x^{j}_{k+1}=x^{j}_{k}-h_{1}{D}_{\mu}F(\mu_{\textbf{x}_{k}},x^{j}_{k})+\sqrt{2\lambda_{1}h_{1}}\xi^{i}_{k}$ (NLA)
for $i=1,...,N$, $k=1,...,K$ and $\xi^{i}_{k}\sim\mathcal{N}(0,I_{d})$, and EM-UNLA (the EM discretization of the NULD with stepsize $h_{2}$ and scaling scalar $\lambda_{2}$) whose update is given by
$x_{k+1}^{j}$ $=x_{k}^{j}+h_{2}v_{k}^{j}$ (EM-NULA)
$v_{k+1}^{j}$ $=(1-\gamma h_{2})v_{k}^{j}-h_{2}{D}_{\mu}F(\mu_{\textbf{x}_{k}},x_{k}^{j})+\sqrt{2\lambda_{2}h_{2}}\xi^{i}_{k}$
[…]
Figure 2: NULA with different number of particles
6 Discussion
This paper (1) improves the convergence guarantees in Chen et al. (2023) with a refined Lyapunov analysis (Theorems 3.1 and 3.2); (2) discretizes the MULD and NULD with a scheme which results in smaller bias than the EM scheme; and (3) presents a novel discretization analysis of MULD and NULD. Our theoretical results suggest the mixing time of MULA and NULA might be superior to that of MLA and NLA, which is further validated by our experiments. We now note several directions for future potential developments. First, it is intriguing to see whether we can improve the number of particle guarantees for NULA to avoid the behavior of convergence instability in the low particle approximation regime. Second, we suspect convergence rates for MULD and NULD can be improved. Importantly, we obtain convergence rates for the MULA and NULA in TV distance, which do not share the same distance with the convergence rates of MULD, NULD, MLA and NLA. We hope to establish our results in the energy gap or KL divergence in the future. What’s more, our techniques on uniformly bounding the iterates of MULD and NULD combined with Assumptions 2.5-2.7 generates an additional $\mathscr{C}_{\textsf{LSI}}$ by Talagrand’s inequality, which is the reason why we do not improve the dependence of $\mathscr{C}_{\textsf{LSI}}$ for MULA and NULA. We hope to explore whether it is possible to weaken those assumptions and refine the analysis of uniformly bounding the iterates to improve the dependence of $\mathscr{C}_{\textsf{LSI}}$ in the mixing time and number of particles of MULA and NULA.
Appendix A Helpful lemmas
Lemma 1.
The solution $(x_{t},v_{t})$ to the discrete-time process (MULA) for $t\in[kh,(k+1)h]$ is
$x_{t}$ $=x_{k}+\frac{1-e^{-\gamma(t-kh)}}{\gamma}v_{k}-\frac{\gamma h-(1-e^{-\gamma(t-kh)})}{\gamma^{2}}D_{\mu}F(\mu_{k}^{x},x_{k})+\mathrm{B}_{k}^{x},$ (19)
$v_{t}$ $=e^{-\gamma(t-kh)}v_{k}-\frac{1-e^{-\gamma(t-kh)}}{\gamma}D_{\mu}F(\mu_{k}^{x},x_{k})+\mathrm{B}_{k}^{v},$
where $(\mathrm{B}_{k}^{x},\mathrm{B}_{k}^{v})\in\mathbb{R}^{2d}$ is independent of $k$ and has the joint distribution
$\left[\begin{array}[]{c}\mathrm{B}_{k}^{x}\\ \mathrm{B}_{k}^{v}\end{array}\right]\sim\mathcal{N}\left(0,\left[\begin{array}[]{cc}\frac{2}{\gamma}\left(h-\frac{2(1-e^{-\gamma(t-kh)})}{\gamma}+\frac{1-e^{-2\gamma(t-kh)}}{2\gamma}\right)&\frac{1}{\gamma}\left(1-2e^{-\gamma(t-kh)}+e^{-2\gamma(t-kh)}\right)\\ &1-e^{-2\gamma(t-kh)}\end{array}\right]\right)$
The solution $(x_{t}^{i},v_{t}^{i})$ to the discrete-time process (7) for $i=1,...,N$ and $t\in[kh,(k+1)h]$ is
$x^{i}_{t}$ $=x^{i}_{k}+\frac{1-e^{-\gamma(t-kh)}}{\gamma}v^{i}_{k}-\frac{\gamma h-(1-e^{-\gamma(t-kh)})}{\gamma^{2}}D_{\mu}F(\mu_{\textbf{x}_{k}},x^{i}_{k})+(\mathrm{B}^{i}_{k})^{x},$ (20)
$v^{i}_{t}$ $=e^{-\gamma(t-kh)}v^{i}_{k}-\frac{1-e^{-\gamma(t-kh)}}{\gamma}D_{\mu}F(\mu_{\textbf{x}_{k}},x^{i}_{k})+(\mathrm{B}^{i}_{k})^{v}.$
where $((\mathrm{B}^{i}_{k})^{x},(\mathrm{B}^{i}_{k})^{v})\in\mathbb{R}^{2d}$ is independent of $i,\,k$ and has the joint distribution
$\left[\begin{array}[]{c}(\mathrm{B}_{k}^{i})^{x}\\ (\mathrm{B}_{k}^{i})^{v}\end{array}\right]\sim\mathcal{N}\left(0,\left[\begin{array}[]{cc}\frac{2}{\gamma}\left(h-\frac{2(1-e^{-\gamma(t-kh)})}{\gamma}+\frac{1-e^{-2\gamma(t-kh)}}{2\gamma}\right)&\frac{1}{\gamma}\left(1-2e^{-\gamma(t-kh)}+e^{-2\gamma(t-kh)}\right)\\ &1-e^{-2\gamma(t-kh)}\end{array}\right]\right)$
Proof.
The proof technique is similar to the proof of Lemmas 10 and 11 proposed in Cheng et al. (2018). ∎
Choosing $t=(k+1)h$ for (20) generates the updates of Algorithm 1.
Lemma 2.
Suppose $D_{\mu}F:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ admits a continuous first variation ${\delta D_{\mu}F}:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Then, $D_{\mu}F$ is $\mathscr{L}$-Lipschitz with respect to $W_{1}$ distance satisfying
$\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x)\|\leq\mathscr{L}W_{1}(\mu,\mu^{\prime})$ (21)
with $\mathscr{L}\triangleq\sup_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\left\|{\delta D_{\mu}F}(\mu,x,x^{\prime})\right\|_{\textsf{\em op}}$
Proof.
By the definition of functional derivative, we have
$\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x)\|\leq\int_{0}^{1}\left\|\int{\delta D_{\mu}F}((1-t)\mu+t\mu^{\prime},x,x^{\prime})(\mu-\mu^{\prime})\mathrm{d}x^{\prime}\right\|\mathrm{d}t$ (22)
By Kantorovich duality, we obtain
$\left\|\int{\delta D_{\mu}F}((1-t)\mu+t\mu^{\prime},x,x^{\prime})(\mu-\mu^{\prime})\mathrm{d}x^{\prime}\right\|\leq\mathscr{L}W_{1}(\mu,\mu^{\prime}).$
Combining with (22), we complete the proof. ∎
Lemma 3 (Mean-field Entropy Sandwich, Chen et al. 2023, Lemma 4.2).
Assume $F$ satisfies Assumptions 2.1-2.3. Then for every $\mu\in\mathcal{P}_{2}(\mathbb{R}^{2d})$ we have
$\textsf{\em KL}(\mu\|\mu_{*})\leq\mathcal{F}(\mu)-\mathcal{F}(\mu_{*})\leq\textsf{\em KL}(\mu\|\hat{\mu})\leq\left(1+\frac{\mathscr{L}}{\mathscr{C}_{\textsf{\em LSI}}}+\frac{\mathscr{L}^{2}}{2\mathscr{C}_{\textsf{\em LSI}}^{2}}\right)\textsf{\em KL}(\mu\|\mu_{*}).$ (23)
Lemma 4 (Particle System’s Entropy Inequality, Chen et al. 2023, Lemma 4.2).
Assume that $F$ satisfies Assumption 2.1 and there exists a measure $\mu_{*}\in\mathcal{P}(\mathbb{R}^{2d})$ that admits the proximal Gibbs distribution $\mu_{*}(x,v)\propto\exp\left(-\frac{\delta F}{\delta\mu}(\mu^{x}_{*},x)-\frac{1}{2}\|v\|^{2}\right)$. Then for all $\mu^{N}\in\mathcal{P}(\mathbb{R}^{2dN})$, we have
$\textsf{\em KL}(\mu^{N}\|\mu_{*}^{\otimes N})\leq\mathcal{F}^{N}(\mu^{N})-N\mathcal{F}(\mu_{*}).$ (24)
Lemma 5 (Information Inequality).
Let $X_{1},...,X_{N}$ be measurable spaces, $\mu$ be a probability on the product space $X=X_{1}\times...\times X_{N}$ and $\nu=\nu^{1}\otimes...\otimes\nu^{N}$ be a $\sigma$-finite measure. Then
$\sum_{i=1}^{N}\textsf{\em KL}(\mu^{i}\|\nu^{i})\leq\textsf{\em KL}(\mu\|\nu).$ (25)
Lemma 6 (Matrix Grönwall’s Inequality, Zhang et al. 2023).
Let $x:\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$, and $c\in\mathbb{R}^{d}$, $A\in\mathbb{R}^{d\times d}$, where $A$ has non-negative entries. Suppose that the following inequality is satisfied componentwise:
$x(t)\leq c+\int_{0}^{t}Ax(s)\mathrm{d}s,\quad\text{for all }t\geq 0.$
Then the following inequality holds where $I_{d}\in\mathbb{R}^{d\times d}$ is the $d$-dimensional identity matrix:
$x(t)\leq\left(AA^{{\dagger}}e^{At}-AA^{{\dagger}}+I_{d}\right)c.$
Lemma 7.
Let $(x_{t},v_{t})_{t\geq 0}$ and $(x_{t}^{i},v_{t}^{i})_{t\geq 0}$ respectively denote the iterates of the MULD and NULD. Assume that $h\lesssim\mathscr{L}^{-1/2}\wedge\gamma^{-1}$. Under Assumption 2.2 and Assumption 2.4, for $t\in[kh,(k+1)h]$, we have
$\sup_{t\in[kh,(k+1)h]}\|x_{t}-x_{kh}\|\leq 2\mathscr{L}h^{2}\|x_{kh}\|+4h\|v_{kh}\|+2\mathscr{L}h^{2}+2\sqrt{2\gamma}h\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|$
$\sup_{t\in[kh,(k+1)h]}\|x^{i}_{t}-x^{i}_{kh}\|\leq 2\mathscr{L}h^{2}\|x^{i}_{kh}\|+4h\|v^{i}_{kh}\|+2\mathscr{L}h^{2}+2\sqrt{2\gamma}h\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}^{i}_{t}-\mathrm{B}^{i}_{kh}\|$
for $i=1,...,N$.
Proof.
We only prove the first relation, and the proof of the second relation is similar.
$\|x_{t}-x_{kh}\|$ $=\left\|\int_{kh}^{t}v_{\tau}\mathrm{d}\tau\right\|\leq h\|v_{kh}\|+\left\|\int_{kh}^{t}v_{\tau}-v_{kh}\mathrm{d}\tau\right\|$
$\leq h\|v_{kh}\|+\left\|\int_{kh}^{t}\int_{0}^{\tau}\gamma v_{\tau^{\prime}}\mathrm{d}\tau^{\prime}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\int_{kh}^{\tau}{D}_{\mu}F(\mu_{\tau^{\prime}}^{x},x_{\tau^{\prime}})\mathrm{d}\tau^{\prime}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\int_{kh}^{\tau}\sqrt{2\gamma}\mathrm{dB}_{\tau^{\prime}}\mathrm{d}\tau\right\|$
$\leq h\|v_{kh}\|+\gamma h\left(h\|v_{kh}\|+\int_{kh}^{t}\|v_{\tau}-v_{kh}\|\mathrm{d}\tau\right)+\left\|\int_{kh}^{t}\int_{kh}^{\tau}{D}_{\mu}F(\mu_{\tau^{\prime}}^{x},x_{\tau^{\prime}})\mathrm{d}\tau^{\prime}\mathrm{d}\tau\right\|$
$\quad+\left\|\int_{kh}^{t}\int_{kh}^{\tau}\sqrt{2\gamma}\mathrm{dB}_{\tau^{\prime}}\mathrm{d}\tau\right\|$
$\leq h\|v_{kh}\|+\gamma h\left(h\|v_{kh}\|+\int_{kh}^{t}\|v_{\tau}-v_{kh}\|\mathrm{d}\tau\right)+\mathscr{L}h\int_{kh}^{t}\left\|x_{\tau}-x_{kh}\right\|\mathrm{d}\tau+\mathscr{L}h^{2}\|x_{kh}\|$
$\quad+\mathscr{L}h^{2}+\sqrt{2\gamma}h\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|$
where the last inequality follows from Assumptions 2.2 and 2.4. Likewise for $V$:
$\|v_{t}-v_{kh}\|$ $=\left\|\int_{kh}^{t}\gamma v_{\tau}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}{D}_{\mu}F(\mu_{\tau}^{x},x_{\tau})\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\sqrt{2\gamma}\mathrm{dB}_{t}\right\|$
$\leq\gamma\left(h\|v_{kh}\|+\int_{kh}^{t}\|v_{\tau}-v_{kh}\|\mathrm{d}\tau\right)+\left\|\int_{kh}^{t}{D}_{\mu}F(\mu_{\tau}^{x},x_{\tau})\mathrm{d}\tau\right\|+\sqrt{2\gamma}\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|$
$\leq\gamma\left(h\|v_{kh}\|+\int_{kh}^{t}\|v_{\tau}-v_{kh}\|\mathrm{d}\tau\right)+\mathscr{L}\int_{kh}^{t}\|x_{\tau}-x_{kh}\|\mathrm{d}\tau+\mathscr{L}h+\mathscr{L}h\|x_{kh}\|$
$\quad+\sqrt{2\gamma}\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|$
where the last inequality follows from Assumptions 2.2 and 2.4. Before applying matrix form of Grönwall’s inequality, let $c=c_{1}+c_{2}$ with $c_{2}=\left[\begin{array}[]{c}h\|v_{kh}\|\\ 0\end{array}\right]$,
$A=\left[\begin{array}[]{cc}\mathscr{L}h&\gamma h\\ \mathscr{L}&\gamma\end{array}\right],\,c_{1}=\left[\begin{array}[]{c}\mathscr{L}h^{2}\|x_{kh}\|+\gamma h^{2}\|v_{kh}\|+\mathscr{L}h^{2}+\sqrt{2\gamma}h\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|\\ \mathscr{L}h\|x_{kh}\|+\gamma h\|v_{kh}\|+\mathscr{L}h+\sqrt{2\gamma}\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|\end{array}\right].$
$c_{1}$ lies in the image space of $A$, and $\exp(A_{t})c_{1}$ also lies in the image space of $A$. For the first component:
$\sup_{t\in[kh,(k+1)h]}\|x_{t}$ $-x_{kh}\|$
$\leq h\exp\left((\mathscr{L}h+\gamma)h\right)(\mathscr{L}h\|x_{kh}\|+\gamma h\|v_{kh}\|+\mathscr{L}h+\sqrt{2\gamma}\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|)$
$\quad+\frac{\mathscr{L}h\exp((\mathscr{L}h+\gamma)h)+\gamma}{\mathscr{L}h+\gamma}h\|v_{kh}\|$
$\leq 2h\left(\mathscr{L}h\|x_{kh}\|+2\|v_{kh}\|+\mathscr{L}h+\sqrt{2\gamma}\sup_{t\in[kh,(k+1)h]}\|\mathrm{B}_{t}-\mathrm{B}_{kh}\|\right)$
where the second inequality comes from choosing $h\lesssim\frac{1}{\mathscr{L}^{1/2}}\land\frac{1}{\gamma}$.
$((AA^{\dagger}(\exp(Ah)-I)+I)c_{2})_{(1)}=\frac{\mathscr{L}h\exp((\mathscr{L}h+\gamma)h)+\gamma}{\mathscr{L}h+\gamma}h\|v_{kh}\|\leq 2h\|v_{kh}\|$
Combining relations above and Lemma 6 completes the proof. ∎
Lemma 8.
Let $(x_{t},v_{t})_{t\geq 0}$ denote the iterates of the MULD with $(x_{0},v_{0})\sim\mu_{0}=\mathcal{N}(0,I_{2d})$. Under Assumption 2.5 and Assumption 2.6, we have
$\mathbb{E}\|(x_{t},v_{t})\|^{2}\lesssim\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{\em LSI}}}$ (26)
Proof.
$\mathbb{E}\|(x_{t},v_{t})\|^{2}=W_{2}^{2}(\mu_{t},\delta_{0})$ $\leq 2W_{2}^{2}(\mu_{t},\mu_{*})+2W_{2}^{2}(\mu_{*},\delta_{0})$
$\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\textsf{KL}(\mu_{t}\|\mu_{*})+2\textbf{m}_{2}^{2}$
$\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))+2\textbf{m}_{2}^{2}$
$\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}(\mathcal{F}(\mu_{0})-\mathcal{F}(\mu_{*}))+2\textbf{m}_{2}^{2}$
$\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\mathcal{F}(\mu_{0})+2\textbf{m}_{2}^{2}$
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 2.3.222Assumption 2.3 states that the proximal Gibbs distribution satisfies the LSI. Note that $\mu_{*}$ also has the form of the proximal Gibbs distribution and thus satisfies LSI. The third inequality follows from Lemma 3. The fourth inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}(\mu_{t})<0$ along the MULD (Proof of Theorem 2.1 in Chen et al. (2023)) and the last inequality follows from the assumption that $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}(\mu)$, we have $\mathcal{F}(\mu_{0})=F(\mu^{x}_{0})+\int\frac{1}{2}\|v\|^{2}\mu_{0}(\mathrm{d}x\mathrm{d}v)+\text{Ent}(\mu_{0})$. Since $(x_{0},v_{0})\sim\mathcal{N}(0,I_{2d})$, we have $\int\frac{1}{2}\|v\|^{2}\mu_{0}(\mathrm{d}x\mathrm{d}v)\lesssim d$ and
$|\text{Ent}(\mu_{0})|$ $=\left|\int\mu_{0}\log\mu_{0}\right|$
$=\frac{d}{2}\log(2\pi)+\frac{1}{2}\mathbb{E}_{\mu_{0}}\|\cdot\|^{2}\lesssim d.$
By Assumption 2.6, we have $F(\mu_{0}^{x})\lesssim\mathscr{L}d$. By Assumption 2.5, we have $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
$\mathbb{E}\|(x_{t},v_{t})\|^{2}$ $\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\mathcal{F}(\mu_{0})+2\textbf{m}_{2}^{2}$
$\lesssim\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}+d$
Lemma 9.
Let ${(x^{i}_{t},v^{i}_{t})_{i=1}^{N}}$ denote the iterates of the NULD with $(x_{0}^{i},v_{0}^{i})\sim\mu_{0}^{i}=\mathcal{N}(0,I_{2d})$ for $i=1,...,N$ and $t\geq 0$. Under Assumption 2.5 and Assumption 2.7, we have
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|(x^{i}_{t},v^{i}_{t})\|^{2}\lesssim\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{\em LSI}}}$ (27)
Proof.
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|(x^{i}_{t},v^{i}_{t})\|^{2}$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{i}_{t},\delta_{0})$
$\leq\frac{2}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{i}_{t},\mu_{*})+2W_{2}^{2}(\mu_{*},\delta_{0})$
$\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mu^{i}_{t}\|\mu_{*})+2\textbf{m}_{2}^{2}$
$\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\frac{1}{N}\textsf{KL}(\mu^{N}_{t}\|\mu^{\otimes N}_{*})+2\textbf{m}_{2}^{2}$
$\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\left(\frac{1}{N}\mathcal{F}^{N}(\mu^{N}_{t})-\mathcal{F}(\mu_{*})\right)+2\textbf{m}_{2}^{2}$
$\leq\frac{2}{N\mathscr{C}_{\textsf{LSI}}}\mathcal{F}^{N}(\mu^{N}_{0})+2\textbf{m}_{2}^{2}$
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 2.3. The third inequality follows from Lemma 5. The fourth inequality follows from Lemma 4 and the last inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}^{N}(\mu^{N}_{t})<0$ along the NULD (Proof of Theorem 2.2 in Chen et al. (2023)) and $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}^{N}(\mu^{N})$, we have $\mathcal{F}^{N}(\mu_{0}^{N})=\int(NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu_{0}^{N})$. Similar to the proof of Lemma 8, since $(\textbf{x},\textbf{v})\sim\mathcal{N}(0,I_{2Nd})$, we have $\int\frac{1}{2}\|\textbf{v}\|^{2}\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim Nd$ and $|\text{Ent}(\mu_{0}^{N})|\lesssim Nd$. By Assumption 2.7 and Assumption 2.5, we also have $\int NF(\mu_{\textbf{x}})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim N\mathscr{L}d$ and $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|(x^{i}_{t},v^{i}_{t})\|^{2}$ $\leq\frac{2}{N\mathscr{C}_{\textsf{LSI}}}\mathcal{F}^{N}(\mu^{N}_{0})+2\textbf{m}_{2}^{2}$
$=\frac{2}{N\mathscr{C}_{\textsf{LSI}}}\left(\int(NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu_{0}^{N})\right)+2\textbf{m}_{2}^{2}$
$\lesssim\frac{1}{N\mathscr{C}_{\textsf{LSI}}}(N\mathscr{L}d+Nd)+d$
$\lesssim\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}+d$
Lemma 10 (Girsanov’s Theorem, (Zhang et al. (2023), Theorem 19)).
Consider stochastic processes $(x_{t})_{t\geq 0}$, $(b_{t}^{\textbf{P}})_{t\geq 0}$, $(b_{t}^{\textbf{Q}})_{t\geq 0}$ adapted to the same filtration, and $\sigma\in\mathbb{R}^{d\times d}$ any constant matrix (possibly degenerate). Let $\textbf{P}_{T}$ and Q be probability measures on the path space $C([0,T];\mathbb{R}^{d})$ such that $(x_{t})_{t\geq 0}$ follows
$\mathrm{d}x_{t}$ $=b_{t}^{\textbf{P}}\mathrm{d}t+\sigma\mathrm{dB}^{\textbf{P}}_{t}\quad\text{under }\textbf{P}_{T},$
$\mathrm{d}x_{t}$ $=b_{t}^{\textbf{Q}}\mathrm{d}t+\sigma\mathrm{dB}^{\textbf{Q}}_{t}\quad\text{under }\textbf{Q}_{T},$
where $\mathrm{B}^{\textbf{P}}$ and $\mathrm{B}^{\textbf{Q}}$ are $\textbf{P}_{T}$-Brownian motion and $\textbf{Q}_{T}$-Brownian motion. Suppose there exists a process $(y_{t})_{t\geq 0}$ such that
$\sigma y_{t}=b_{t}^{\textbf{P}}-b_{t}^{\textbf{Q}},$
and
$\mathbb{E}_{\mathbf{Q}_{T}}\exp\left(\frac{1}{2}\int_{0}^{T}\left\|y_{t}\right\|^{2}\mathrm{d}t\right)<\infty.$
If we define $\sigma^{{\dagger}}$ as the Moore-Penrose pseudo-inverse of $\sigma$, then we have
$\frac{\mathrm{d}\mathbf{P}_{T}}{\mathrm{d}\mathbf{Q}_{T}}=\exp\left(\int_{0}^{T}\langle\sigma_{t}^{{\dagger}}(b_{t}^{\textbf{P}_{T}}-b_{t}^{\textbf{Q}_{T}}),\mathrm{dB}^{\textbf{Q}_{T}}_{t}\rangle-\frac{1}{2}\int_{0}^{T}\|\sigma_{t}^{{\dagger}}(b_{t}^{\textbf{P}_{T}}-b_{t}^{\textbf{Q}_{T}})\|^{2}\mathrm{d}t\right)$
Besides, $(\tilde{\mathrm{B}}_{t})_{t\in[0,T]}$ defined by $\mathrm{d}\tilde{\mathrm{B}}_{t}\triangleq\mathrm{dB}_{t}+\sigma_{t}^{{\dagger}}(b_{t}^{Y}-b_{t}^{X})$ is a $\mathbf{P}_{T}$-Brownian motion.
Appendix B Verification of assumptions
B.1 Verification of Assumption 2.2
$W_{1}$ smoothness has been verified for training mean-field neural networks in Chen et al. (2022). Thus we only verify $W_{1}$ smoothness for examples of density estimation via MMD minimization and KSD minimization. Lemma 2 provides sufficient conditions for $W_{1}$ smoothness. In particular, we have
$\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x^{\prime})\|\leq\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x)\|+\|D_{\mu}F(\mu^{\prime},x)-D_{\mu}F(\mu^{\prime},x^{\prime})\|$ (28)
Suzuki et al. (2023) verify that $\|D_{\mu}F(\mu^{\prime},x)-D_{\mu}F(\mu^{\prime},x^{\prime})\|\leq\mathscr{L}\|x-x^{\prime}\|$ for three examples mentioned above. Thus it suffices to verify (21) for the last two examples.
MMD minimization
We now prove that objective (15) satisfies Assumption 2.2 with Gaussian RBF kernel and $r(x)=\|x\|^{2}/2$. We choose $\sigma^{\prime}$ in Gaussian RBF kernel $k$ to be $\sigma$ for brevity. With the specific choice of $r$, we reformulate (15) as
$F(\mu)=\hat{\mathcal{M}}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}.$ (29)
According to the definition of ^$\hat{\mathcal{M}}(\mu)$ in Section 4, the intrinsic derivative of $F(\mu)$ is
$D_{\mu}F(\mu,x)$ $=D_{\mu}\hat{\mathcal{M}}(\mu,x)+\frac{\lambda^{\prime}}{2}\|x\|^{2}$
$={2\iiint\nabla_{x}p(x;z)p(x^{\prime};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}\mathrm{d}\mu(x^{\prime})}-{\frac{2}{n}\sum_{i=1}^{n}\int\nabla_{x}p(x;z)k(z,z_{i})\mathrm{d}z}+\frac{\lambda^{\prime}}{2}\|x\|^{2}$
We only need to prove $D_{\mu}\hat{\mathcal{M}}(\mu,x)$ is smooth. The first variation of $D_{\mu}\hat{\mathcal{M}}(\mu,x)$ is
${\delta D_{\mu}\hat{\mathcal{M}}}(\mu,x,x^{\prime})$ $=2\iint\nabla_{x}p(x;z)p(x^{\prime};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}$
$=-\frac{2}{(2\pi\sigma^{2})^{d}\sigma^{2}}\iint(x-z)\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}\right)\exp\left(-\frac{\|x^{\prime}-z^{\prime}\|^{2}+\|z-z^{\prime}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z\mathrm{d}z^{\prime}$
Thus we have
$\left\|{\delta D_{\mu}\hat{\mathcal{M}}}(\mu,x,x^{\prime})\right\|$ $\leq\frac{2}{(2\pi\sigma^{2})^{d}\sigma^{2}}\iint\|x-z\|\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}\right)\exp\left(-\frac{\|x^{\prime}-z^{\prime}\|^{2}+\|z-z^{\prime}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z\mathrm{d}z^{\prime}$
$\leq\frac{2}{(2\pi\sigma^{2})^{d}\sigma}\iint\exp\left(-\frac{\|x^{\prime}-z^{\prime}\|^{2}+\|z-z^{\prime}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z\mathrm{d}z^{\prime}=\frac{2}{\sigma}$
According to Lemma 2 and (28), $F(\mu)$ defined in (29) satisfies Assumption 2.2.
KSD minimization
We now prove that objective (17) satisfies Assumption 2.2 with kernel
$k(x,x^{\prime})=\exp\left(-\frac{\|x\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x^{\prime}\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x-x^{\prime}\|^{2}}{2\sigma_{2}^{2}}\right)$ (30)
and $r(x)=\|x\|^{2}/2$. We also assume the score function of $\mu_{*}$ satisfies (18). Under this assumption on score function and with this choice of kernel, Suzuki et al. (2023) show in their Appendix A that the Stein kernel $u_{\mu_{*}}$ satisfies $\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\max\{|u_{\mu_{*}}|,\|\nabla_{x}u_{\mu_{*}}\|,\|\nabla_{x}^{2}u_{\mu_{*}}\|_{\textsf{op}}\}\leq\mathscr{L}$. With the specific choice of $r(x)=\|x\|^{2}/2$, we reformulate (17) as
$F(\mu)=\textsf{KSD}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}.$ (31)
Similarly, we only need to verify that $\textsf{KSD}(\mu)$ is smooth with respect to $W_{1}$ distance. The intrinsic derivative of $\textsf{KSD}(\mu)$ is
$D_{\mu}\textsf{KSD}(\mu,x)=\int\nabla_{x}u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x^{\prime}).$
The first variation of $D_{\mu}\textsf{KSD}(\mu,x)$ is
$\delta D_{\mu}\textsf{KSD}(\mu,x,x^{\prime})=\nabla_{x}u_{\mu_{*}}(x,x^{\prime})$
The following relation implies Assumption 2.2 by Lemma 2.
$\|\delta D_{\mu}\textsf{KSD}(\mu,x,x^{\prime})\|=\|\nabla_{x}u_{\mu_{*}}(x,x^{\prime})\|\leq\mathscr{L}$
B.2 Verification of Assumption 2.5
Training mean-field neural networks
Denote $\hat{\mu}(x,v)=\hat{\mu}^{x}(x)\otimes\mathcal{N}(0,I_{d})$ where $\hat{\mu}^{x}(x)\propto\exp\left(-\frac{\delta F}{\delta\mu}(\mu^{x},x)\right)$. Since the second moment of $\mathcal{N}(0,I_{d})$ is $O(d)$, it suffices to ensure $\mathbb{E}_{x\sim\hat{\mu}^{x}(x)}\|x\|^{2}=O(d)$. Consider the empirical risk minimization problem in the mean-field neural networks setting with training data $(a_{i},b_{i})_{i=1}^{n}$:
$F(\mu)=\frac{1}{n}\sum_{i=1}^{n}l(h(\mu;a_{i}),b_{i})+\lambda^{\prime}\mathbb{E}_{x\sim\mu}[r(x)].$ (32)
  • We will prove that Assumption 2.5 holds if $r(x)=\|x\|^{2}/2$, $|h(x;a)|\leq\sqrt{\mathscr{L}}$ (such activation functions include tanh and sigmoid) and $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$ (such loss functions include logistic loss, Huber loss and log-cosh loss) or $l$ is quadratic. Since $r(x)=\|x\|^{2}/2$, we obtain
    $\frac{\delta F}{\delta\mu}(\mu,x)$ $=\frac{1}{n}\sum_{i=1}^{n}\left[\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})\right]+\frac{\lambda^{\prime}}{2}\|x\|^{2}$
    Consider the case where $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$. Since $|h(x;a)|\leq\sqrt{\mathscr{L}}$, we have $|\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})|\leq\mathscr{L}$. Let $\hat{\mu}^{x}(x)={\exp\left(-\frac{\delta F}{\delta\mu}(\mu,x)\right)}/{Z}$ where $Z=\int\exp\left(-\frac{\delta F}{\delta\mu}(\mu,x)\right)\mathrm{d}x$, and we have
    $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}$ $=\frac{1}{Z}\int\|x\|^{2}\exp\left(-\frac{1}{n}\sum_{i=1}^{n}\left[\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})\right]-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x\triangleq\frac{Z^{\prime}}{Z}$ (33)
    Now we bound $Z^{\prime}$ and $Z$ respectively.
    $Z^{\prime}$ $\leq\int\|x\|^{2}\exp\left(\mathscr{L}-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x\lesssim\frac{\exp(\mathscr{L})d}{\lambda^{\prime}},$
    $Z$ $\geq\int\exp\left(-\mathscr{L}-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x=\exp(-\mathscr{L})\left(\frac{2\pi}{\lambda^{\prime}}\right)^{d/2}$
    Choose $\lambda^{\prime}$ satisfying $\lambda^{\prime}\lesssim\frac{(2\pi)^{\frac{d}{d-2}}}{\exp\left(\frac{4\mathscr{L}}{d-2}\right)}$, and we have $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=\frac{Z^{\prime}}{Z}\lesssim\frac{\exp(2\mathscr{L})}{\lambda^{\prime}\left(\frac{2\pi}{\lambda^{\prime}}\right)^{d/2}}d\lesssim d.$ Consider the case where $l$ is quadratic. $|h(\mu;a_{i})|=|\int h(x;a_{i})\mu(\mathrm{d}x)|\leq\int|h(x;a_{i})|\mu(\mathrm{d}x)\leq\sqrt{\mathscr{L}}$, thus we have $|\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})|=|(h(\mu;a_{i})-b_{i})h(x;a_{i})|\leq\mathscr{L}+|b_{i}|\sqrt{\mathscr{L}}$. We can scale the training data to ensure $\max_{i=1}^{n}|b_{i}|\leq\sqrt{\mathscr{L}}$, and we obtain $|\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})|\leq 2\mathscr{L}$. The remaining proof keeps the same with $\lambda^{\prime}\lesssim\frac{(2\pi)^{\frac{d}{d-2}}}{\exp\left(\frac{8\mathscr{L}}{d-2}\right)}$.
  • We will prove that Assumption 2.5 holds if $r(x)=\|x\|^{2}/2$, $|h(x;a)|\leq\sqrt{\mathscr{L}}(1+\|x\|)$ (such activation functions include ReLU, GeLU, Softplus, SiLU) and $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$. Under these conditions, we have $|\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})|\leq\mathscr{L}(1+\|x\|)$. Then, based on (33), we obtain
    $Z^{\prime}\leq\int\|x\|^{2}\exp\left(\mathscr{L}(1+\|x\|)-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x$ $\leq\exp(\mathscr{L})\int\|x\|^{2}\exp\left(\frac{3\mathscr{L}^{2}}{2\lambda^{\prime}}-\frac{\lambda^{\prime}}{3}\|x\|^{2}\right)\mathrm{d}x$
    $\lesssim\exp\left(\mathscr{L}+\frac{3\mathscr{L}^{2}}{2\lambda^{\prime}}\right)\frac{d}{\lambda^{\prime}}.$
    We also have
    $Z\geq\int\exp\left(-\mathscr{L}(1+\|x\|)-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x$ $\gtrsim\exp(\mathscr{L})\int\exp\left(-\frac{\mathscr{L}^{2}}{\lambda^{\prime}}-\frac{3\lambda^{\prime}}{4}\|x\|^{2}\right)\mathrm{d}x$
    $=\exp\left(\mathscr{L}-\frac{\mathscr{L}^{2}}{\lambda^{\prime}}\right)\left(\frac{4\pi}{3\lambda^{\prime}}\right)^{d/2}$
    Combining the upper bound of $Z^{\prime}$ and the lower bound of $Z$, if $d\gtrsim\frac{\mathscr{L}^{2}}{\lambda^{\prime}}$, we obtain
    $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=\frac{Z^{\prime}}{Z}\lesssim\exp\left(\frac{5\mathscr{L}^{2}}{2\lambda^{\prime}}\right)\frac{d}{\lambda^{\prime}}\left(\frac{3\lambda^{\prime}}{4\pi}\right)^{d/2}\leq\exp\left(\frac{5\mathscr{L}^{2}}{2\lambda^{\prime}}\right)\left(\frac{3}{4\pi}\right)^{d/2}d\lesssim d.$
    Note that $d\gtrsim\frac{\mathscr{L}^{2}}{\lambda^{\prime}}$ is possible for large-scale problems.
MMD minimization
We now prove that objective (15) satisfies Assumption 2.5 with Gaussian RBF kernel and $r(x)=\|x\|^{2}/2$. We choose $\sigma^{\prime}$ in Gaussian RBF kernel $k$ to be $\sigma$ for brevity. With the specific choice of $r$, we reformulate (15) as
$F(\mu)=\hat{\mathcal{M}}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}.$ (34)
According to the definition of ^$\hat{\mathcal{M}}(\mu)$ in Section 4, the functional derivative of ^$\hat{\mathcal{M}}(\mu)$ is
$\frac{\delta\hat{\mathcal{M}}(\mu)}{\delta\mu}(x)=\underbrace{2\iiint p(x;z)p(x^{\prime};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}\mathrm{d}\mu(x^{\prime})}_{\textsf{P}}-\underbrace{\frac{2}{n}\sum_{i=1}^{n}\int p(x;z)k(z,z_{i})\mathrm{d}z}_{\textsf{Q}}$ (35)
Next we bound each part of $\frac{\delta\hat{\mathcal{M}}(\mu)}{\delta\mu}(x)$. For P, we have
$\frac{1}{2}\textsf{P}$ $=\frac{1}{(2\pi\sigma^{2})^{d}}\iiint\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}-\frac{\|x^{\prime}-z^{\prime}\|^{2}}{2\sigma^{2}}-\frac{\|z-z^{\prime}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z\mathrm{d}z^{\prime}\mathrm{d}\mu(x^{\prime})$
$=\frac{(\pi\sigma^{2})^{\frac{d}{2}}}{(2\pi\sigma^{2})^{d}}\iint\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{6\sigma^{2}}-\frac{3\|z^{\prime}-\frac{2}{3}x^{\prime}-\frac{1}{3}x\|^{2}}{4\sigma^{2}}\right)\mathrm{d}z^{\prime}\mathrm{d}\mu(x^{\prime})$
$=\left(\frac{1}{\sqrt{3}}\right)^{d}\int\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{6\sigma^{2}}\right)\mathrm{d}\mu(x^{\prime})\leq\left(\frac{1}{\sqrt{3}}\right)^{d}$
where the last inequality follows from the relation $\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{6\sigma^{2}}\right)\leq 1$. For $\mathsf{Q}$, we have
$\frac{1}{2}\mathsf{Q}$ $=\frac{1}{(2\pi\sigma^{2})^{\frac{d}{2}}}\frac{1}{n}\sum_{i=1}^{n}\int\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}-\frac{\|z-z_{i}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z$
$=\frac{1}{(2\pi\sigma^{2})^{\frac{d}{2}}}\frac{1}{n}\sum_{i=1}^{n}\exp\left(-\frac{\|x\|^{2}+\|z_{i}\|^{2}}{2\sigma^{2}}+\frac{\|z_{i}+x\|^{2}}{4\sigma^{2}}\right)\ \int\exp\left(-\frac{\|z-\frac{1}{2}z_{i}-\frac{1}{2}x\|^{2}}{\sigma^{2}}\right)\mathrm{d}z$
$=\left(\frac{1}{\sqrt{2}}\right)^{d}\frac{1}{n}\sum_{i=1}^{n}\exp\left(-\frac{\|x\|^{2}+\|z_{i}\|^{2}}{2\sigma^{2}}+\frac{\|z_{i}+x\|^{2}}{4\sigma^{2}}\right)\leq\left(\frac{1}{\sqrt{2}}\right)^{d}$
where the last inequality follows from the relation $\|z_{i}+x\|^{2}\leq 2\|z_{i}\|^{2}+2\|x\|^{2}$. Note that $\mathsf{P}\geq 0$ and $\textsf{Q}\geq 0$. Combining the bound of P and Q, we obtain the bound of $\frac{\delta\hat{\mathcal{M}}(\mu)}{\delta\mu}(x)$ as follows:
$-\sqrt{2}\leq-2\left(\frac{1}{\sqrt{2}}\right)^{d}\leq\frac{\delta\hat{\mathcal{M}}(\mu)}{\delta\mu}(x)=\textsf{P}-\textsf{Q}\leq 2\left(\frac{1}{\sqrt{3}}\right)^{d}\leq\sqrt{3}$ (36)
Let $\hat{\mu}^{x}(x)={\exp\left(-\frac{\delta F}{\delta\mu}(\mu,x)\right)}/{Z}$ where $Z=\int\exp\left(-\frac{\delta F}{\delta\mu}(\mu,x)\right)\mathrm{d}x$, and we have
$\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}$ $=\frac{1}{Z}\int\|x\|^{2}\exp\left(-\frac{\delta\hat{\mathcal{M}}(\mu)}{\delta\mu}(x)-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x\triangleq\frac{Z^{\prime}}{Z}$ (37)
Now we bound $Z^{\prime}$ and $Z$ respectively.
$Z^{\prime}$ $\leq\int\|x\|^{2}\exp\left(\sqrt{2}-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x\lesssim\frac{\exp(\sqrt{2})d}{\lambda^{\prime}},$
$Z$ $\geq\int\exp\left(-\sqrt{3}-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x=\exp(-\sqrt{3})\left(\frac{2\pi}{\lambda^{\prime}}\right)^{d/2}$
Thus we have $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=\frac{Z^{\prime}}{Z}\lesssim\frac{\exp(\sqrt{2}+\sqrt{3})\lambda^{\prime\frac{d}{2}-1}}{(2\pi)^{\frac{d}{2}}}d\lesssim d$ for $\lambda^{\prime}\lesssim 2\pi$.
KSD minimization
Assume the score function $s_{\mu_{*}}$ satisfies (18) and choose the kernel $k$ to be (30), and the Stein kernel $u_{\mu_{*}}$ satisfies $\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\max\{|u_{\mu_{*}}|,\|\nabla_{x}u_{\mu_{*}}\|,\|\nabla_{x}^{2}u_{\mu_{*}}\|_{\textsf{op}}\}\leq\mathscr{L}$ (Suzuki et al., 2023). We now prove the following objective
$F(\mu)=\textsf{KSD}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}$ (38)
satisfies Assumption 2.5, with $\textsf{KSD}(\mu)$ defined in (16). The functional derivative of $\textsf{KSD}(\mu)$ is
$\frac{\delta\textsf{KSD}}{\delta\mu}(\mu,x)=\int u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x^{\prime}).$
The functional derivative is bounded as
$\left|\frac{\delta\textsf{KSD}}{\delta\mu}(\mu,x)\right|\leq\int|u_{\mu_{*}}(x,x^{\prime})|\mathrm{d}\mu(x^{\prime})\leq\mathscr{L}.$
Let $\hat{\mu}^{x}(x)={\exp\left(-\frac{\delta F}{\delta\mu}(\mu,x)\right)}/{Z}$ where $Z=\int\exp\left(-\frac{\delta F}{\delta\mu}(\mu,x)\right)\mathrm{d}x$, and we have
$\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}$ $=\frac{1}{Z}\int\|x\|^{2}\exp\left(-\frac{\delta\textsf{KSD}}{\delta\mu}(\mu,x)-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x\triangleq\frac{Z^{\prime}}{Z}$ (39)
Now we bound $Z^{\prime}$ and $Z$ respectively.
$Z^{\prime}$ $\leq\int\|x\|^{2}\exp\left(\mathscr{L}-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x\lesssim\frac{\exp(\mathscr{L})d}{\lambda^{\prime}},$
$Z$ $\geq\int\exp\left(-\mathscr{L}-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x=\exp(-\mathscr{L})\left(\frac{2\pi}{\lambda^{\prime}}\right)^{d/2}$
Thus we have $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=\frac{Z^{\prime}}{Z}\lesssim\frac{\exp(2\mathscr{L})d\lambda^{\prime\frac{d}{2}-1}}{(2\pi)^{\frac{d}{2}}}\lesssim d$ for $\lambda^{\prime}\lesssim(2\pi)^{\frac{d}{d-2}}\exp\left(-\frac{4\mathscr{L}}{d-2}\right)$.
B.3 Verification of Assumption 2.6
Training mean-field neural networks
Consider the same objective in (14) with $r(x)=\|x\|^{2}/2$ and $\mu_{0}=\mathcal{N}(0,I_{d})$:
$F(\mu)=\frac{1}{n}\sum_{i=1}^{n}l(h(\mu;a_{i}),b_{i})+\lambda^{\prime}\mathbb{E}_{x\sim\mu}[r(x)].$
  • If $l$ is $\sqrt{\mathscr{L}}$-Lipschitz, we have $|l(h(\mu;a),b)|\leq\sqrt{\mathscr{L}}|h(\mu;a)-b|$. If $|h(x;a)|\leq\sqrt{\mathscr{L}}$, we have $|h(\mu;a)|\leq\sqrt{\mathscr{L}}$ and thus $F(\mu_{0})\lesssim\sqrt{\mathscr{L}}(\sqrt{\mathscr{L}}+\max_{i=1}^{n}|b_{i}|)+d$. We can normalize the data samples to ensure $\max_{i=1}^{n}|b_{i}|\lesssim d\land\sqrt{\mathscr{L}}$. Thus $F(\mu_{0})=O(\mathscr{L}+d)$.
  • If $|h(x;a)|\leq\sqrt{\mathscr{L}}(1+\|x\|)$, we have $|h(\mu_{0};a)|\leq\sqrt{\mathscr{L}}\int(1+\|x\|)\mu_{0}(\mathrm{d}x)\lesssim\sqrt{\mathscr{L}}d^{1/2}$. If $l$ is $\sqrt{\mathscr{L}}$-Lipschitz, we have $|l(h(\mu_{0};a_{i}),b_{i})|\leq\sqrt{\mathscr{L}}|h(\mu_{0};a_{i})-b_{i}|\lesssim\mathscr{L}d^{1/2}+\sqrt{\mathscr{L}}\max_{i=1}^{n}|b_{i}|$. We can normalize the data samples to ensure $\max_{i=1}^{n}|b_{i}|\lesssim d\land\sqrt{\mathscr{L}}$. Thus we have $F(\mu_{0})=O(\mathscr{L}d+d)$.
MMD minimization
Consider the same objective in (15) with Gaussian RBF kernel ($\sigma^{\prime}=\sigma$), $r(x)=\|x\|^{2}/2$ and $\mu_{0}=\mathcal{N}(0,I_{d})$:
$F(\mu)=\hat{\mathcal{M}}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2},$ (40)
where
^$\hat{\mathcal{M}}(\mu)$ $=\iiint p(x;z)p(x^{\prime};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}\mathrm{d}(\mu\times\mu)(x,x^{\prime})-2\int\left(\frac{1}{n}\sum_{i=1}^{n}\int p(x;z)k(z,z_{i})\mathrm{d}z\right)\mathrm{d}\mu(x)$
$=\frac{1}{{3}^{d/2}}\int\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{6\sigma^{2}}\right)\mathrm{d}(\mu\times\mu)(x,x^{\prime})-\frac{2}{2^{d/2}}\frac{1}{n}\sum_{i=1}^{n}\int\exp\left(-\frac{\|x-z_{i}\|^{2}}{4\sigma^{2}}\right)\mathrm{d}\mu(x)$
$\leq\frac{1}{{3}^{d/2}}\int\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{6\sigma^{2}}\right)\mathrm{d}(\mu\times\mu)(x,x^{\prime})\leq\frac{1}{{3}^{d/2}}\leq\mathscr{L}$
Thus $F(\mu_{0})=\hat{\mathcal{M}}(\mu_{0})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu_{0}}\|x\|^{2}\lesssim\mathscr{L}+d,$ which satisfies Assumption 2.6.
KSD minimization
Consider the same objective in (17) with $r(x)=\|x\|^{2}/2$ and $\mathcal{N}(0,I_{d})$:
$F(\mu)=\textsf{KSD}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}.$
If we choose kernel $k(x,x^{\prime})=\exp\left(-\frac{\|x\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x^{\prime}\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x-x^{\prime}\|^{2}}{2\sigma_{2}^{2}}\right)$ and assume the score function of $\mu_{*}$ satisfies $\max\{\|\nabla\log\mu_{*}(x)\|,\|\nabla^{\otimes 2}\log\mu_{*}(x)\|_{\textsf{op}},\|\nabla^{\otimes 3}\log\mu_{*}(x)\|_{\textsf{op}}\}\leq\mathscr{L}(1+\|x\|)$, then the Stein kernel $u_{\mu_{*}}$ satisfies $\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\max\{|u_{\mu_{*}}|,\|\nabla_{x}u_{\mu_{*}}\|,\|\nabla_{x}^{2}u_{\mu_{*}}\|_{\textsf{op}}\}\leq\mathscr{L}$ according to the statement of Appendix A in Suzuki et al. (2023). We have
$F(\mu_{0})$ $=\textsf{KSD}(\mu_{0})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu_{0}}\|x\|^{2}$
$=\iint u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu_{0}}\|x\|^{2}$
$\lesssim\mathscr{L}+d,$
which satisfies Assumption 2.6.
B.4 Verification of Assumption 2.7
Training mean-field neural networks
Similar to examples of training mean-field neural networks above, we initialize $\mu^{N}_{0}=\mathcal{N}^{\otimes N}(0,I_{d})$ and choose $r(x)=\|x\|^{2}/2$ for the following objective.
$\mathbb{E}_{\textbf{x}\sim\mu^{N}}F(\mu_{\textbf{x}})\triangleq\mathbb{E}_{\textbf{x}\sim\mu^{N}}\frac{1}{n}\sum_{i=1}^{n}\left[l\left(\frac{1}{N}\sum_{s=1}^{N}h(x^{s};a_{i}),b_{i}\right)\right]+\frac{\lambda^{\prime}}{2}\mathbb{E}_{\textbf{x}\sim\mu^{N}}\frac{1}{N}\sum_{s=1}^{N}\left[\|x^{s}\|^{2}\right],$
where $\textbf{x}=(x^{1},...,x^{N})$, $x^{i}\sim\mu^{i}$ for $i=1,...,N$ and $\mu^{N}=\otimes_{i=1}^{N}\mu^{i}=\text{Law}(x^{1},...,x^{N})$.
  • If $|h(x;a)|\leq\sqrt{\mathscr{L}}$ and $l$ is $\sqrt{\mathscr{L}}$-Lipschitz, and $\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\frac{1}{n}\sum_{i=1}^{n}\left[l\left(\frac{1}{N}\sum_{i=1}^{N}h(x_{0}^{i};a_{i}),b_{i}\right)\right]\lesssim\sqrt{\mathscr{L}}(\sqrt{\mathscr{L}}+\max_{i=1}^{n}|b_{i}|)$ and thus $\mathbb{E}_{\mu_{0}^{N}}F(\mu_{\textbf{x}_{0}})\lesssim\mathscr{L}+\sqrt{\mathscr{L}}\max_{i=1}^{n}|b_{i}|+d$. We can normalize the data samples to ensure $\max_{i=1}^{n}|b_{i}|\lesssim d\land\sqrt{\mathscr{L}}$. Thus we have $\mathbb{E}_{\mu_{0}^{N}}F(\mu_{\textbf{x}_{0}})=O(\mathscr{L}+d)$.
  • If $|h(x;a)|\leq\sqrt{\mathscr{L}}(1+\|x\|)$ and $l$ is $\sqrt{\mathscr{L}}$-Lipschitz, $\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\frac{1}{n}\sum_{i=1}^{n}\left[l\left(\frac{1}{N}\sum_{s=1}^{N}h(x_{0}^{s};a_{i}),b_{i}\right)\right]\leq\sqrt{\mathscr{L}}\left(\sqrt{\mathscr{L}}\frac{1}{N}\sum_{s=1}^{N}(1+\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\|x_{0}^{s}\|)+\max_{i=1}^{n}|b_{i}|\right)\lesssim\mathscr{L}d^{1/2}+\sqrt{\mathscr{L}}\max_{i=1}^{n}|b_{i}|$. We can normalize the data samples to ensure $\max_{i=1}^{n}|b_{i}|\lesssim d\land\sqrt{\mathscr{L}}$. Thus we have $\mathbb{E}_{\mu_{0}^{N}}F(\mu_{\textbf{x}_{0}})=O(\mathscr{L}d+d)$
MMD minimization
Now we verify Assumption 2.7 for the example of density estimation. We consider the N-particle approximation of the objective (40) with $r(x)=\|x\|^{2}/2$ and initialization $\mu^{N}_{0}=\mathcal{N}^{\otimes N}(0,I_{d})$.
$\mathbb{E}_{\mu^{N}}$$\hat{\mathcal{M}}(\mu_{\textbf{x},\textbf{y}})$
$\triangleq\mathbb{E}_{\textbf{x},\textbf{y}\sim\mu^{N}}\left[\frac{1}{N^{2}}\sum_{s=1}^{N}\sum_{t=1}^{N}\iint p(x^{s};z)p(y^{t};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}-\frac{2}{nN}\sum_{i=1}^{n}\sum_{s=1}^{N}\int p(x^{s};z)k(z,z_{i})\mathrm{d}z\right]$
$\leq\mathbb{E}_{\textbf{x},\textbf{y}\sim\mu^{N}}\left[\frac{1}{N^{2}}\sum_{s=1}^{N}\sum_{t=1}^{N}\iint p(x^{s};z)p(y^{t};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}\right]$
$=\left(\frac{1}{\sqrt{3}}\right)^{d}\mathbb{E}_{\textbf{x}\sim\mu^{N}}\left[\frac{1}{N^{2}}\sum_{s=1}^{N}\sum_{t=1}^{N}\exp\left(-\frac{\|x^{s}-y^{t}\|^{2}}{6\sigma^{2}}\right)\right]\leq\left(\frac{1}{\sqrt{3}}\right)^{d}\leq\mathscr{L}$
where $\textbf{x}=(x^{1},...,x^{N})$ and $\textbf{y}=(y^{1},...,y^{N})$. Thus we can upper bound $\mathbb{E}_{\textbf{x}_{0},\textbf{y}_{0}\sim\mu^{N}_{0}}F(\mu_{\textbf{x}_{0},\textbf{y}_{0}})$ as follows:
$\mathbb{E}_{\textbf{x}_{0},\textbf{y}_{0}\sim\mu^{N}_{0}}F(\mu_{\textbf{x}_{0},\textbf{y}_{0}})=\mathbb{E}_{\textbf{x}_{0},\textbf{y}_{0}\sim\mu^{N}_{0}}\hat{\mathcal{M}}(\mu_{\textbf{x},\textbf{y}})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\frac{1}{N}\sum_{s=1}^{N}\left[\|x_{0}^{s}\|^{2}\right]\lesssim\mathscr{L}+d$
which satisfies Assumption 2.7.
KSD minimization
Similar to the verification of Assumption 2.6 above, we have the following relation for $\mu_{0}^{N}=\mathcal{N}(0,I_{Nd})$ under the same assumptions on the score function and kernel:
$\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}}F(\mu_{\textbf{x}_{0}})$ $=\textsf{KSD}(\mu_{\textbf{x}_{0}})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\frac{1}{N}\sum_{s=1}^{N}\left[\|x_{0}^{s}\|^{2}\right]$
$=\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}}\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}u_{\mu_{*}}(x_{0}^{i},x_{0}^{j})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\frac{1}{N}\sum_{s=1}^{N}\left[\|x_{0}^{s}\|^{2}\right]\lesssim\mathscr{L}+d,$
which satisfies Assumption 2.7.
Appendix C Continuous-time results
In this section, we give the explicit rate of Theorem 2.1 and Theorem 2.2 proposed by Chen et al. (2023) with a specific choice of parameters and then provide the detailed proof of Theorem 3.1 and Theorem 3.2 by reparameterizing $\gamma$.
C.1 Proof of Theorem 3.1
Our proof is directly adapted from Theorem 2.1 in Chen et al. (2023) using hypocoercivity in Villani (2009). Chen et al. (2023) prove the Lyapunov functional
$\mathcal{E}(\mu_{t})=\mathcal{F}(\mu_{t})+\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu}_{t})$ (41)
is decaying along the MULD with $S=\left(\begin{array}[]{cc}c&b\\ b&a\end{array}\right)\otimes I_{d}$ and $\gamma=1$. Let $A_{t}=\nabla_{v}$, $B_{t}=v\cdot\nabla_{x}-{D}_{\mu}F(\mu_{t}^{x},x)\cdot\nabla_{v}$, $C_{t}=[A_{t},B_{t}]=A_{t}B_{t}-B_{t}A_{t}=\nabla_{x}$ and $Y_{t}=(\|A_{t}u_{t}\|,\|A_{t}^{2}u_{t}\|,\|C_{t}u_{t}\|,\|C_{t}A_{t}u_{t}\|)^{\mathsf{T}}$ where $u_{t}=\log\frac{\mu_{t}}{\hat{\mu}_{t}}$. More specifically, Chen et al. (2023) prove that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}(\mu_{t})\leq-Y_{t}^{\mathsf{T}}\mathcal{K}Y_{t},$ (42)
where
$\mathcal{K}=\left(\begin{array}[]{cccc}1+2a-4\mathscr{L}b&-2b&-2a-2\mathscr{L}c&0\\ 0&2a&-2\mathscr{L}c&-4b\\ 0&0&2b&0\\ 0&0&0&2c\end{array}\right).$
The choice of $a,\,b,\,c$ should satisfies $ac>b^{2}$ and $K\succ 0$. If we choose $a=c=2\mathscr{L}$ and $b=1$, the smallest eigenvalue of $\mathcal{K}$ is $\lambda_{\textsf{min}}(\mathcal{K})=1$, and thus we have
$\frac{\mathrm{d}}{\mathrm{d}t}(\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*}))$ $\leq-(\|A_{t}u_{t}\|^{2}+\|A_{t}^{2}u_{t}\|^{2}+\|C_{t}u_{t}\|^{2}+\|C_{t}A_{t}u_{t}\|^{2})$
$\leq-(\|A_{t}u_{t}\|^{2}+\|C_{t}u_{t}\|^{2})=-\frac{1}{2}\mathsf{FI}(\mu_{t}\|\hat{\mu}_{t})-\frac{1}{2}\mathsf{FI}(\mu_{t}\|\hat{\mu}_{t})$
$\leq-\mathscr{C}_{\textsf{LSI}}\textsf{KL}(\mu_{t}\|\hat{\mu}_{t})-\frac{1}{2\lambda_{\textsf{max}}(S)}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu_{t}})$
$\leq-\mathscr{C}_{\textsf{LSI}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{4\mathscr{L}+2}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu_{t}})$
$\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{6\mathscr{L}}(\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*}))$
Applying Grönwall’s inequality, we obtain
$\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*})\leq\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*})\leq(\mathcal{E}(\mu_{0})-\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{6\mathscr{L}}t\right).$ (43)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}(\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*}))$ $\leq-\frac{2}{\sqrt{\mathscr{L}}}(\|A_{t}u_{t}\|^{2}+\|A_{t}^{2}u_{t}\|^{2}+\|C_{t}u_{t}\|^{2}+\|C_{t}A_{t}u_{t}\|^{2})$
$\leq-\frac{2}{\sqrt{\mathscr{L}}}(\|A_{t}u_{t}\|^{2}+\|C_{t}u_{t}\|^{2})=-\frac{1}{\sqrt{\mathscr{L}}}\mathsf{FI}(\mu_{t}\|\hat{\mu}_{t})-\frac{1}{\sqrt{\mathscr{L}}}\mathsf{FI}(\mu_{t}\|\hat{\mu}_{t})$
$\leq-\frac{2\mathscr{C}_{\textsf{LSI}}}{\sqrt{\mathscr{L}}}\textsf{KL}(\mu_{t}\|\hat{\mu}_{t})-\frac{1}{\lambda_{\textsf{max}}(S)\sqrt{\mathscr{L}}}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu}_{t})$
$\leq-\frac{2\mathscr{C}_{\textsf{LSI}}}{\sqrt{\mathscr{L}}}(\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*}))-\frac{1}{3\sqrt{\mathscr{L}}}\mathsf{FI}_{S}(\mu_{t}\|\hat{\mu}_{t})$
$\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{3\sqrt{\mathscr{L}}}(\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*}))$
where the fourth inequality follows from $\lambda_{\textsf{max}}(S)=\frac{\frac{1}{\mathscr{L}}+2+\sqrt{\frac{1}{\mathscr{L}^{2}}+4}}{2}\leq\frac{1}{\mathscr{L}}+2\leq 3.$ Applying Grönwall’s inequality, we obtain
$\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*})\leq\mathcal{E}(\mu_{t})-\mathcal{E}(\mu_{*})\leq(\mathcal{E}(\mu_{0})-\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{3\sqrt{\mathscr{L}}}t\right),$ (44)
which completes the proof of Theorem 3.1. Eq. 44 exhibits a faster rate than the rate of Eq. 43.
C.2 Proof of Theorem 3.2
Our proof is directly adapted from Theorem 2.2 in Chen et al. (2023) using hypocoercivity in Villani (2009). Chen et al. (2023) prove that the Lyapunov functional
$\mathcal{E}^{N}(\mu_{t}^{N})=\mathcal{F}^{N}(\mu_{t}^{N})+\textsf{FI}^{N}_{S}(\mu_{t}^{N}\|\mu_{*}^{N})$ (45)
is decaying along the NULD with $S=\left(\begin{array}[]{cc}c&b\\ b&a\end{array}\right)\otimes I_{d}$ and $\gamma=1$. Let $u_{t}^{N}=\log\frac{\mu_{t}^{N}}{\mu^{N}_{*}}$ and
$Y_{t}^{N}=(\|\nabla_{\textbf{v}}u_{t}^{N}\|,\|\nabla_{\textbf{v}}^{2}u_{t}^{N}\|,\|\nabla_{\textbf{x}}u_{t}^{N}\|,\|\nabla_{\textbf{x}}\nabla_{\textbf{v}}u_{t}^{N}\|)^{\mathsf{T}}.$
Chen et al. (2023) prove that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}^{N}(\mu_{t}^{N})\leq-(Y_{t}^{N})^{\mathsf{T}}\mathcal{K}Y_{t}^{N}$ (46)
where
$\mathcal{K}=\left(\begin{array}[]{cccc}1+2a-4\mathscr{L}b&-2b&-2a&0\\ 0&2a&-4\mathscr{L}c&-4b\\ 0&0&2b&0\\ 0&0&0&2c\end{array}\right).$
The choice of $a,\,b,\,c$ should satisfies $ac>b^{2}$ and $K\succ 0$. If we choose $a=c=2\mathscr{L}$ and $b=1$, the smallest eigenvalue of $K$ is $\lambda_{\textsf{min}}(\mathcal{K})=1$, and thus we have
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}^{N}(\mu_{t}^{N})$ $\leq-(\|\nabla_{\textbf{v}}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{v}}^{2}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{x}}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{x}}\nabla_{\textbf{v}}u_{t}^{N}\|^{2})$ (47)
$\leq-(\|\nabla_{\textbf{v}}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{x}}u_{t}^{N}\|^{2})=-\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})$
Since $\mu_{*}^{N}$ does not satisfy the uniform LSI, we can not utilize the same technique to upper bound $-\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})$. Chen et al. (2022) and Chen et al. (2023) obtain the lower bound of the relative Fisher information $\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})$ using other technique to circumvent the uniform LSI of $\mu_{*}^{N}$. We will directly provide the conclusion instead of providing many details about that technique in this paper, and we refer our readers to Chen et al. (2022, 2023) for the precise proof. Chen et al. (2023) propose that
$\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})$ $=\frac{1}{2}\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})+\frac{1}{2}\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})$
$\geq\frac{1}{2}\left[2(1-\varepsilon)\mathscr{C}_{\textsf{LSI}}-\frac{\mathscr{L}}{N}\left(16+12(\varepsilon^{-1}-1)\frac{\mathscr{L}}{\mathscr{C}_{\textsf{LSI}}}\right)\right](\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*}))$
$\quad+\frac{1}{2}\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})-\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3(\varepsilon^{-1}-1)\mathscr{L})$
for $\varepsilon\in(0,1)$. If we choose $\varepsilon=1/2$ and $N\geq\frac{32\mathscr{L}}{\mathscr{C}_{\textsf{LSI}}}+\frac{24\mathscr{L}^{2}}{\mathscr{C}_{\textsf{LSI}}^{2}}$, we have
$\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})$ $\geq\frac{\mathscr{C}_{\textsf{LSI}}}{4}(\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*}))+\frac{1}{2\lambda_{\textsf{max}}(S)}\textsf{FI}_{S}(\mu_{t}^{N}\|\mu_{*}^{N})-\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})$
$\geq\frac{\mathscr{C}_{\textsf{LSI}}}{4}(\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*}))+\frac{1}{6\mathscr{L}}\textsf{FI}_{S}(\mu_{t}^{N}\|\mu_{*}^{N})-\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})$
$\geq\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}(\mathcal{E}^{N}(\mu^{N}_{t})-N\mathcal{E}(\mu_{*}))-\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})$
Combining (47) with the lower bound of Fisher information above, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}(\mathcal{E}^{N}(\mu_{t}^{N})-N\mathcal{E}(\mu_{*}))\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}(\mathcal{E}^{N}(\mu^{N}_{t})-N\mathcal{E}(\mu_{*}))+\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})$
Applying Grönwall’s inequality, we obtain
$\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*})$ $\leq\mathcal{E}^{N}(\mu_{t}^{N})-N\mathcal{E}(\mu_{*})$ (48)
$\leq(\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}t\right)+\frac{\mathscr{L}\mathrm{d}t}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}t\right)$
$\leq(\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}t\right)+\frac{120\mathscr{L}^{2}d}{\mathscr{C}_{\textsf{LSI}}}+\frac{72\mathscr{L}^{3}d}{\mathscr{C}_{\textsf{LSI}}^{2}}$
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}(\mathcal{E}^{N}(\mu_{t}^{N})-N\mathcal{E}(\mu_{*}))$ $\leq-\frac{2}{\sqrt{\mathscr{L}}}(\|\nabla_{\textbf{v}}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{v}}^{2}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{x}}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{x}}\nabla_{\textbf{v}}u_{t}^{N}\|^{2})$
$\leq-\frac{2}{\sqrt{\mathscr{L}}}(\|\nabla_{\textbf{v}}u_{t}^{N}\|^{2}+\|\nabla_{\textbf{x}}u_{t}^{N}\|^{2})=-\frac{2}{\sqrt{\mathscr{L}}}\textsf{FI}(\mu_{t}^{N}\|\mu_{*}^{N})$
$\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{2\sqrt{\mathscr{L}}}(\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*}))-\frac{1}{\lambda_{\textsf{max}}(S)\sqrt{\mathscr{L}}}\textsf{FI}_{S}(\mu_{t}^{N}\|\mu_{*}^{N})$
$\quad+\frac{2\sqrt{\mathscr{L}}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})$
$\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{2\sqrt{\mathscr{L}}}(\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*}))-\frac{1}{3\sqrt{\mathscr{L}}}\textsf{FI}_{S}(\mu_{t}^{N}\|\mu_{*}^{N})+\frac{2\sqrt{\mathscr{L}}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})$
$\leq-\frac{\mathscr{C}_{\textsf{LSI}}}{6\sqrt{\mathscr{L}}}(\mathcal{E}^{N}(\mu^{N}_{t})-N\mathcal{E}(\mu_{*}))+\frac{2\sqrt{\mathscr{L}}d}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})$
Applying Grönwall’s inequality, we obtain
$\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*})\leq\mathcal{E}^{N}(\mu_{t}^{N})-N\mathcal{E}(\mu_{*})\leq\mathcal{E}_{0}^{N}\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{6\sqrt{\mathscr{L}}}t\right)+\frac{60\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}+\frac{36\mathscr{L}^{2}d}{\mathscr{C}_{\textsf{LSI}}^{2}}$ (49)
where $\mathcal{E}_{0}^{N}\triangleq\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}).$ This completes the proof of Theorem 3.2. The convergence rate exhibited in Eq. 49 is faster and incurs a smaller bias than the rate exhibited in Eq. 48.
Appendix D Discretization analysis
In this section, we provide the proof of Theorem 3.3 and Theorem 3.4 establishing the global convergence of the discrete-time-space processes. Our discretization analysis is unified for the MULA and NULA.
D.1 Proof of Theorem 3.3
Suppose $\textbf{Q}_{Nh}$ is the joint law of the MULD for $t\in[0,Nh]$ and $\textbf{P}_{Nh}$ is the joint law of the MULA for $t\in[kh,(k+1)h]$ and $k=0,1,...,K-1$. Applying Girsanov’s theorem (Lemma 10), we have
$\textsf{KL}(\mathbf{Q}_{Kh}\|\mathbf{P}_{Kh})$ $=\mathbb{E}_{\mathbf{Q}_{Kh}}\log\frac{\mathrm{d}\mathbf{Q}_{Kh}}{\mathrm{d}\mathbf{P}_{Kh}}$
$=\mathbb{E}_{\mathbf{Q}_{Kh}}\sum_{k=0}^{K-1}\left(-\frac{1}{\sqrt{2\gamma}}\int_{kh}^{(k+1)h}\left\langle\left(\begin{array}[]{c}0\\ {D}_{\mu}F(\mu_{t}^{x},x_{t})-{D}_{\mu}F(\mu_{kh}^{x},x_{kh})\end{array}\right),\mathrm{dB}_{t}\right\rangle\right.$
$\left.\quad+\frac{1}{4\gamma}\int_{kh}^{(k+1)h}\left\|{D}_{\mu}F(\mu_{t}^{x},x_{t})-{D}_{\mu}F(\mu_{kh}^{x},x_{kh})\right\|^{2}\mathrm{d}t\right)$
$=\frac{1}{4\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{Q}_{Kh}}\left\|{D}_{\mu}F(\mu_{t}^{x},x_{t})-{D}_{\mu}F(\mu_{kh}^{x},x_{kh})\right\|^{2}\mathrm{d}t$
And we obtain
$\textsf{KL}(\mathbf{Q}_{Kh}\|\mathbf{P}_{Kh})$ $=\frac{1}{4\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{Q}_{Kh}}\left\|{D}_{\mu}F(\mu_{t}^{x},x_{t})-{D}_{\mu}F(\mu_{kh}^{x},x_{kh})\right\|^{2}\mathrm{d}t$
$\leq\frac{{\mathscr{L}}^{2}}{2\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{t}-x_{kh}\|^{2}+W_{1}^{2}(\mu_{t}^{x},\mu_{kh}^{x})\mathrm{d}t$
$\leq\frac{\mathscr{L}^{2}}{2\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{t}-x_{kh}\|^{2}+\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{t}-x_{kh}\|^{2}\mathrm{d}t$
$=\frac{\mathscr{L}^{2}}{\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{t}-x_{kh}\|^{2}\mathrm{d}t$
$\leq\frac{16\mathscr{L}^{4}h^{5}}{\gamma}\sum_{k=0}^{K-1}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}}{\gamma}\sum_{k=0}^{K-1}\mathbb{E}_{\mathbf{Q}_{Kh}}\|v_{kh}\|^{2}+\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}+{32\mathscr{L}^{2}h^{4}Kd}$
where the first inequality follows from Assumption 2.2 and the last inequality follows from Lemma 7 and the inequality $\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\right)^{2}\leq\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}$:
$\mathbb{E}_{\textbf{Q}_{Kh}}\|x_{t}-x_{kh}\|^{2}\leq 16\mathscr{L}^{2}h^{4}\mathbb{E}_{\textbf{Q}_{Kh}}\|x_{kh}\|^{2}+64h^{2}\mathbb{E}_{\textbf{Q}_{Kh}}\|v_{kh}\|^{2}+16\mathscr{L}^{2}h^{4}+32\gamma h^{3}d$
Combined with Lemma 8 and $\gamma=\sqrt{\mathscr{L}}$, the discretization error is upper bounded as follows:
$\textsf{KL}(\mathbf{Q}_{Kh}\|\mathbf{p}_{Kh})$ $\leq\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}K}{\gamma}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}_{Kh}}\|v_{kh}\|^{2}$
$\quad+\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}+{32\mathscr{L}^{2}h^{4}Kd}$
$\lesssim\frac{\mathscr{L}^{9/2}h^{5}Kd}{\mathscr{C}_{\textsf{LSI}}}+\frac{\mathscr{L}^{5/2}h^{3}Kd}{\mathscr{C}_{\textsf{LSI}}}+{\mathscr{L}^{7/2}h^{5}K}+{\mathscr{L}^{2}h^{4}Kd}$
$=\frac{\mathscr{L}^{9/2}h^{4}Td}{\mathscr{C}_{\textsf{LSI}}}+\frac{\mathscr{L}^{5/2}h^{2}Td}{\mathscr{C}_{\textsf{LSI}}}+{\mathscr{L}^{7/2}h^{4}T}+{\mathscr{L}^{2}h^{3}Td}$
where $T=Kh$. By Lemma 3 and Theorem 3.1, we obtain
$\textsf{KL}(\mu_{t}\|\mu_{*})\leq\mathcal{F}(\mu_{t})-\mathcal{F}(\mu_{*})\leq(\mathcal{E}(\mu_{0})-\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{3\sqrt{\mathscr{L}}}t\right)$ (50)
Combining with (50), we upper bound the TV distance between $\bar{\mu}_{K}$, the probability measure of MULA at $Kh$ and $\mu_{*}$, the limiting distribution of MULD as follows:
$\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}$ $\leq\|\bar{\mu}_{K}-\mu_{Kh}\|_{\textsf{TV}}+\|\mu_{Kh}-\mu_{*}\|_{\textsf{TV}}$
$=\|\mu_{Kh}-\bar{\mu}_{K}\|_{\textsf{TV}}+\|\mu_{Kh}-\mu_{*}\|_{\textsf{TV}}$
$\lesssim\sqrt{\textsf{KL}(\mu_{Kh}\|\bar{\mu}_{K})}+\sqrt{\textsf{KL}(\mu_{Kh}\|\mu_{*})}$
$\lesssim\sqrt{\textsf{KL}(\mathbf{Q}_{Kh}\|\mathbf{p}_{Kh})}+\sqrt{\textsf{KL}(\mu_{Kh}\|\mu_{*})}$
$\lesssim\frac{\mathscr{L}^{9/4}h^{2}T^{1/2}d^{1/2}}{\mathscr{C}^{1/2}_{\textsf{LSI}}}+\frac{\mathscr{L}^{5/4}hT^{1/2}d^{1/2}}{\mathscr{C}^{1/2}_{\textsf{LSI}}}+{\mathscr{L}^{7/4}h^{2}T^{1/2}}+{\mathscr{L}h^{3/2}T^{1/2}d^{1/2}}$
$\quad+(\mathcal{E}(\mu_{0})-\mathcal{E}(\mu_{*}))^{1/2}\exp\left(-{\mathscr{C}_{\textsf{LSI}}T}/{6\sqrt{\mathscr{L}}}\right)$
where the first inequality follows from the triangle inequality of TV distance; the second inequality follows from Pinsker’s inequality, and the fourth inequality follows from the data processing inequality. In order to ensure $\|\mu_{Kh}-\mu_{*}\|_{\textsf{TV}}\leq\frac{1}{2}\epsilon$, it suffices to choose $T=Kh=\widetilde{\Theta}\left(\frac{\sqrt{\mathscr{L}}}{\mathscr{C}_{\textsf{LSI}}}\right)$. In order to ensure $\|\bar{\mu}_{K}-\mu_{Kh}\|_{\textsf{TV}}\leq\frac{1}{2}\epsilon$, it suffices to choose the stepsize
$h={\Theta}\left(\frac{\mathscr{C}^{1/2}_{\textsf{LSI}}\epsilon}{\mathscr{L}^{5/4}T^{1/2}d^{1/2}}\right)=\widetilde{\Theta}\left(\frac{\mathscr{C}_{\textsf{LSI}}\epsilon}{\mathscr{L}^{3/2}d^{1/2}}\right),$ (51)
and the mixing time
$K=\frac{T}{h}=\widetilde{\Theta}\left(\frac{\mathscr{L}^{2}d^{1/2}}{\mathscr{C}_{\textsf{LSI}}^{2}\epsilon}\right).$ (52)
The choice of $T,\,h,\,K$ above ensures $\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}\leq\epsilon$.
D.2 Proof of Theorem 3.4
Suppose $Q_{Nh}^{i}$ is the joint law of the NULD for the $i$-th particle and $t\in[0,Kh]$; $P_{Nh}^{i}$ is the joint law of the NULA for the $i$-th particle. Applying Girsanov’s theorem (Lemma 10), we have
$\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mathbf{Q}^{i}_{Kh}\|\mathbf{P}^{i}_{Kh})$ $=\frac{1}{4\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\left\|{D}_{\mu}F({\mu_{\textbf{x}_{t}}},x^{i}_{t})-{D}_{\mu}F(\mu_{\textbf{x}_{kh}},x^{i}_{kh})\right\|^{2}\mathrm{d}t$
$\leq\frac{\mathscr{L}^{2}}{2\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{t}-x^{i}_{kh}\|^{2}+W_{1}^{2}({\mu_{\textbf{x}_{t}}},\mu_{\textbf{x}_{kh}})\mathrm{d}t$
$\leq\frac{\mathscr{L}^{2}}{\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{t}-x^{i}_{kh}\|^{2}\mathrm{d}t$
$\leq\frac{16\mathscr{L}^{4}h^{5}}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|v_{kh}\|^{2}$
$\quad+\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}+{32\mathscr{L}^{2}h^{4}Kd}$
where the first inequality follows from Assumption 2.2 and the last inequality follows from Lemma 7 and the inequality $\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\right)^{2}\leq\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}$:
$\mathbb{E}_{\textbf{Q}^{i}_{Kh}}\|x^{i}_{t}-x^{i}_{kh}\|^{2}\leq 16\mathscr{L}^{2}h^{4}\mathbb{E}_{\textbf{Q}^{i}_{Kh}}\|x^{i}_{kh}\|^{2}+64h^{2}\mathbb{E}_{\textbf{Q}^{i}_{Kh}}\|v^{i}_{kh}\|^{2}+16\mathscr{L}^{2}h^{4}+32\gamma h^{3}d$
for $t\in[kh,(k+1)h]$ and $k=0,1,...,K-1$. Combining Lemma 9 and $\gamma=\sqrt{\mathscr{L}}$, the discretization error is upper bounded as follows:
$\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mathbf{Q}^{i}_{Kh}\|\mathbf{P}^{i}_{Kh})$ $\leq\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}K}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|v_{kh}\|^{2}$
$\quad+\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}+{32\mathscr{L}^{2}h^{4}Kd}$
$\lesssim\frac{\mathscr{L}^{9/2}h^{5}Kd}{\mathscr{C}_{\textsf{LSI}}}+\frac{\mathscr{L}^{5/2}h^{3}Kd}{\mathscr{C}_{\textsf{LSI}}}+{\mathscr{L}^{7/2}h^{5}K}+{\mathscr{L}^{2}h^{4}Kd}$
$=\frac{\mathscr{L}^{9/2}h^{4}Td}{\mathscr{C}_{\textsf{LSI}}}+\frac{\mathscr{L}^{5/2}h^{2}Td}{\mathscr{C}_{\textsf{LSI}}}+{\mathscr{L}^{7/2}h^{4}T}+{\mathscr{L}^{2}h^{3}Td}$
where $T=Kh$. By Lemma 4 and Theorem 3.2, we obtain
$\frac{1}{N}\textsf{KL}(\mu_{T}^{N}\|\mu_{*}^{\otimes N})\leq\frac{1}{N}\mathcal{F}^{N}(\mu_{T}^{N})-\mathcal{F}(\mu_{*})\leq\frac{\mathcal{E}_{0}^{N}}{N}\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{6\sqrt{\mathscr{L}}}T\right)+\frac{60\mathscr{L}d}{N\mathscr{C}_{\textsf{LSI}}}+\frac{36\mathscr{L}^{2}d}{N\mathscr{C}_{\textsf{LSI}}^{2}},$ (53)
where $\mathcal{E}_{0}^{N}\triangleq\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}).$ Combining with (53), we upper bound the averaged TV distance between $\bar{\mu}_{K}^{i}$ and $\mu_{*}$ over $N$ particles as follows:
$\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}_{K}^{i}-{\mu}_{*}\|_{\textsf{TV}}$ $\leq\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}_{K}^{i}-\mu_{Kh}^{i}\|_{\textsf{TV}}+\frac{1}{N}\sum_{i=1}^{N}\|\mu_{Kh}^{i}-\mu_{*}\|_{\textsf{TV}}$
$=\frac{1}{N}\sum_{i=1}^{N}\|\mu_{Kh}^{i}-\bar{\mu}_{K}^{i}\|_{\textsf{TV}}+\frac{1}{N}\sum_{i=1}^{N}\|\mu_{Kh}^{i}-\mu_{*}\|_{\textsf{TV}}$
$\lesssim\frac{1}{N}\sum_{i=1}^{N}\sqrt{\textsf{KL}({\mu}_{Kh}^{i}\|\bar{\mu}_{K}^{i})}+\frac{1}{N}\sum_{i=1}^{N}\sqrt{\textsf{KL}(\mu_{Kh}^{i}\|\mu_{*})}$
$\lesssim\sqrt{\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}({\mu}_{Kh}^{i}\|\bar{\mu}_{K}^{i})}+\sqrt{\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mu_{Kh}^{i}\|\mu_{*})}$
$\leq\sqrt{\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mathbf{Q}_{Kh}^{i}\|\mathbf{P}_{Kh}^{i})}+\sqrt{\frac{1}{N}\textsf{KL}(\mu_{Kh}^{N}\|\mu^{\otimes N}_{*})}$
$\leq\sqrt{\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mathbf{Q}_{Kh}^{i}\|\mathbf{P}_{Kh}^{i})}+\sqrt{\frac{1}{N}\mathcal{F}^{N}(\mu_{Kh}^{N})-\mathcal{F}(\mu_{*})}$
$\lesssim\frac{\mathscr{L}^{9/4}h^{2}T^{1/2}d^{1/2}}{\mathscr{C}^{1/2}_{\textsf{LSI}}}+\frac{\mathscr{L}^{5/4}hT^{1/2}d^{1/2}}{\mathscr{C}^{1/2}_{\textsf{LSI}}}+{\mathscr{L}^{7/4}h^{2}T^{1/2}}+{\mathscr{L}h^{3/2}T^{1/2}d^{1/2}}$
$\quad+\left(\frac{1}{N}\mathcal{E}^{N}(\mu_{0}^{N})-\mathcal{E}(\mu_{*})\right)^{1/2}\exp\left(-{\mathscr{C}_{\textsf{LSI}}T}/{12\sqrt{\mathscr{L}}}\right)+\frac{\mathscr{L}^{1/2}d^{1/2}}{N^{1/2}\mathscr{C}^{1/2}_{\textsf{LSI}}}+\frac{\mathscr{L}d^{1/2}}{N^{1/2}\mathscr{C}_{\textsf{LSI}}}$
where the first inequality follows from the triangle inequality of TV distance; the second inequality follows from Pinsker’s inequality; the third inequality follows from Jensen’s inequality; the fourth inequality follows from data processing inequality and the information inequality (Lemma 5) and the fifth inequality follows from Lemma 4. In order to ensure $\frac{1}{N}\sum_{i=1}^{N}\|\mu^{i}_{Kh}-\mu_{*}\|_{\textsf{TV}}\leq\frac{1}{2}\epsilon$, it suffices to choose $T=Kh=\widetilde{\Theta}\left(\frac{\sqrt{\mathscr{L}}}{\mathscr{C}_{\textsf{LSI}}}\right)$. In order to ensure $\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}^{i}_{K}-\mu^{i}_{Kh}\|_{\textsf{TV}}\leq\frac{1}{2}\epsilon$, it suffices to choose the stepsize
$h={\Theta}\left(\frac{\mathscr{C}^{1/2}_{\textsf{LSI}}\epsilon}{\mathscr{L}^{5/4}T^{1/2}d^{1/2}}\right)=\widetilde{\Theta}\left(\frac{\mathscr{C}_{\textsf{LSI}}\epsilon}{\mathscr{L}^{3/2}d^{1/2}}\right),$ (54)
the mixing time
$K=\frac{T}{h}=\widetilde{\Theta}\left(\frac{\mathscr{L}^{2}d^{1/2}}{\mathscr{C}_{\textsf{LSI}}^{2}\epsilon}\right),$ (55)
and the number of particles
$N=\Theta\left(\frac{\mathscr{L}^{2}d}{\mathscr{C}_{\textsf{LSI}}^{2}\epsilon^{2}}\right).$ (56)
The choice of $T,\,h,\,K,\,N$ above ensures $\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}^{i}_{K}-\mu_{*}\|_{\textsf{TV}}\leq\epsilon$.
Appendix E Experimental settings
In our experiment, we use a mean-field two-layer neural network to approximate the Gaussian function,
$f(z)=\exp\left(-\frac{\|z-m\|^{2}}{2d}\right).$
We uniformly draw $m\sim\mathcal{N}(0,I_{d})$ and $100$ points $\{z_{i}\}_{i=1}^{100}\sim\mathcal{N}(0,I_{d})$ with $d=10^{3}$ and calculate the corresponding labels $\{f(z_{i})\}_{i=1}^{100}$. In this section, we give the actual updates of the methods involved in our experiment and provide the precise value of parameters in Table 2. The update of the NULA is given by
$\qquad x^{j}_{k+1}$ $=x^{j}_{k}+\varphi_{0}\,v^{j}_{k}-\varphi_{1}\,{D}_{\mu}F(\mu_{\textbf{x}_{k}},x^{j}_{k})+\eta\xi^{x}_{k},$
$\qquad v^{j}_{k+1}$ $=\varphi_{2}\,v^{j}_{k}-\varphi_{3}\,{D}_{\mu}F(\mu_{\textbf{x}_{k}},x^{j}_{k})+\eta\xi^{v}_{k}.$
for $j=1,...,N$. The update of EM-NULA is given by
$\qquad\ \quad\quad x_{k+1}^{j}$ $=x_{k}^{j}+h_{2}\,v_{k}^{j},$
$\qquad\ \quad\quad v_{k+1}^{j}$ $=(1-h_{3})v_{k}^{j}-h_{2}\,{D}_{\mu}F(\mu_{\textbf{x}_{k}},x^{j}_{k})+\sqrt{2\lambda_{2}h_{2}}\xi_{k}.$
for $j=1,...,N$. The update of the NLA is given by
$\ \ x^{j}_{k+1}=x^{j}_{k}-h_{1}\,{D}_{\mu}F(\mu_{\textbf{x}_{k}},x^{j}_{k})+\sqrt{2\lambda_{1}h_{1}}\xi_{k}.$
for $j=1,...,N$.
Parameters $\varphi_{0}$ $\varphi_{1}$ $\varphi_{2}$ $\varphi_{3}$ $\eta$ $h_{1}$ $h_{2}$ $h_{3}$ $\lambda_{1}$ $\lambda_{2}$
Value $10^{-4}$ $0.02$ $0.99$ $0.02$ $10^{-3}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-4}$ $10^{-4}$
Table 2: Choice of hyperparameters.
Appendix F Methods for comparisons
In this section, we review the convergence result of MLA in Nitanda et al. (2022) and NLA in Suzuki et al. (2023), which consider problem (1) in more specific settings. Nitanda et al. (2022) suppose $F(\mu)=\mathbb{E}_{(a,b)\sim\mathcal{D}}\left[l(h(\mu;a),b)\right]+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}$ whereas Suzuki et al. (2023) suppose $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$. While our convergence results are established in TV distance, we consider more general settings compared with the previous two. Since the problem setting in Nitanda et al. (2022) is only for training neural networks, we perform convergence analysis of the MLA in Suzuki et al. (2023)’s setting to make a comparison with our results. Define the free energy
$\mathcal{L}(\mu)=F(\mu)+\text{Ent}(\mu),$ (57)
where $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Let $\bar{\mu}_{k}$ denotes the law of $k$-th iterate of the MLA and $\mu_{*}$ denotes the minimizer of (57), and Nitanda et al. (2022) obtain the following results in Theorem 2:
$\mathcal{L}(\bar{\mu}_{k})-\mathcal{L}(\mu_{*})\leq\exp(-\mathscr{C}_{\textsf{LSI}}hk)(\mathcal{L}(\bar{\mu}_{0})-\mathcal{L}(\mu_{*}))+\frac{\delta_{h}}{2\mathscr{C}_{\textsf{LSI}}},$ (58)
where $\delta_{hk}\triangleq\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k+1},x_{k+1})-{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}$ and $\mathbb{E}$ is taken under the joint law of $\bar{\mu}_{k+1}$ and $\bar{\mu}_{k}$. Now we bound $\delta_{hk}$ uniformly in $k$ with a different method from the one in Nitanda et al. (2022). We do not need to specify $F$ to be the objective of training nerual networks. Since $F$ is $\mathscr{L}$-smooth [^3] We inherit the weaker smoothness assumption in Suzuki et al. (2023) with respect to $W_{2}$ distance. and satisfies Assumption 2.4, we obtain
$\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k+1},x_{k+1})-{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}$ $\leq 2\mathscr{L}^{2}\mathbb{E}(\|x_{k+1}-x_{k}\|^{2}+W_{2}^{2}(\bar{\mu}_{k+1},\bar{\mu}_{k}))$
$\leq 4\mathscr{L}^{2}\mathbb{E}\|x_{k+1}-x_{k}\|^{2}$
$=4\mathscr{L}^{2}\mathbb{E}\|-h{D}_{\mu}F(\bar{\mu}_{k},x_{k})+\sqrt{2h}\xi\|^{2}$
$\leq 4\mathscr{L}^{2}h^{2}\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}+8\mathscr{L}^{2}hd$
$\leq 8\mathscr{L}^{4}h^{2}(1+\mathbb{E}\|x_{k}\|^{2})+8\mathscr{L}^{2}hd$
We refer to Lemma 1 in Suzuki et al. (2023) to uniformly bound $\mathbb{E}\|x_{k}\|^{2}$. Before applying Lemma 1, we translate some constants in Suzuki et al. (2023) into our constants systems. Suzuki et al. (2023) assumes that $\|{D}_{\mu}U(\mu,x)\|\leq R$, $\lambda_{1}I_{d}\preceq\nabla^{2}r(x)\preceq\lambda_{2}I_{d}$. We let $R=\mathscr{L}$ and $\lambda_{2}=\mathscr{L}$ (since this specification matches our Assumption 2.4). We prove Lemma 1 proposed by Suzuki et al. (2023) in the mean-field setting without particle approximation. But we also assume the decomposition $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ with $\|{D}_{\mu}U(\mu,x)\|\leq\mathscr{L}$ and $\lambda_{1}I_{d}\preceq\nabla^{2}r\preceq\mathscr{L}I_{d}$. Given the update of the MLA, if $h\leq\frac{\lambda_{1}}{2\mathscr{L}^{2}}$, we have
$\mathbb{E}\|x_{k+1}\|^{2}$ $=\mathbb{E}\|x_{k}\|^{2}+h^{2}\mathbb{E}\|{D}_{\mu}F(\mu_{k},x_{k})\|^{2}+2hd-2h\mathbb{E}\left\langle x_{k},{D}_{\mu}U(\mu_{k},x_{k})+\nabla r(x_{k})\right\rangle$
$\leq\mathbb{E}\|x_{k}\|^{2}+\mathscr{L}^{2}h^{2}(1+\mathbb{E}\|x_{k}\|^{2})+2hd+2h\mathscr{L}\mathbb{E}\|x_{k}\|-2h\lambda_{1}\mathbb{E}\|x_{k}\|^{2}$
$\leq(1-\lambda_{1}h)\mathbb{E}\|x_{k}\|^{2}+\mathscr{L}^{2}h^{2}+2hd+\frac{2\mathscr{L}^{2}h}{\lambda_{1}}$
Recursively, we obtain
$\mathbb{E}\|x_{k}\|^{2}\leq(1-\lambda_{1}h)^{k}\mathbb{E}\|x_{0}\|^{2}+\frac{\mathscr{L}^{2}h+2d}{\lambda_{1}}+\frac{2\mathscr{L}^{2}}{\lambda_{1}^{2}}\leq\mathbb{E}\|x_{0}\|^{2}+\frac{\mathscr{L}^{2}h+2d}{\lambda_{1}}+\frac{2\mathscr{L}^{2}}{\lambda_{1}^{2}}.$ (59)
If $x_{0}\sim\mathcal{N}(0,I_{d})$, $\mathbb{E}\|x_{0}\|^{2}=O(d)$. Thus (59) implies $\mathbb{E}\|x_{k}\|^{2}=O(\mathscr{L}^{2}d)$. Plugging into the inequality above, we obtain
$\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k+1},x_{k+1})-{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}\lesssim\mathscr{L}^{6}h^{2}d+\mathscr{L}^{2}hd.$ (60)
Applying Lemma 3 and pinsker’s inequality, we obtain
$\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}\lesssim\sqrt{\textsf{KL}(\bar{\mu}_{K}\|\mu_{*})}$ $\leq\sqrt{\mathcal{L}(\bar{\mu}_{K})-\mathcal{L}(\mu_{*})}$
$\lesssim\exp(-\mathscr{C}_{\textsf{LSI}}hK/2)(\mathcal{L}(\bar{\mu}_{0})-\mathcal{L}(\mu_{*}))^{1/2}+\frac{\mathscr{L}^{3}hd^{1/2}}{\mathscr{C}_{\textsf{LSI}}^{1/2}}+\frac{\mathscr{L}h^{1/2}d^{1/2}}{\mathscr{C}_{\textsf{LSI}}^{1/2}}$
In order to ensure $\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}\leq\epsilon$, it suffices to choose
$h=\Theta\left(\frac{\mathscr{C}_{\textsf{LSI}}\epsilon^{2}}{\mathscr{L}^{3}d}\right),\quad K=\widetilde{\Theta}\left(\frac{\mathscr{L}^{3}d}{\mathscr{C}_{\textsf{LSI}}^{2}\epsilon^{2}}\right).$ (61)
Now we translate the convergence results in Suzuki et al. (2023). Define the free energy of the particle system:
$\mathcal{L}^{N}(\mu^{N})=N\mathbb{E}_{\textbf{x}\sim\mu^{N}}F(\mu_{\textbf{x}})+\text{Ent}(\mu^{N}),$ (62)
where $\mu_{\textbf{x}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}$. Similar to the analysis above, Theorem 2 in Suzuki et al. (2023) implies the TV-convergence of the NLA, given by
$\frac{1}{N}\sum_{i=1}^{N}\|\bar{\mu}_{K}^{i}-{\mu}_{*}\|_{\textsf{TV}}\lesssim\sqrt{\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\bar{\mu}_{K}^{i}\|\mu_{*})}$ $\leq\sqrt{\frac{1}{N}\mathcal{L}^{N}(\mu_{K}^{N})-\mathcal{L}(\mu_{*})}$
$\lesssim\exp\left(-\mathscr{C}_{\textsf{LSI}}hK/4\right)+h^{1/2}K^{1/2}(\mathscr{L}^{3}hd^{1/2}+\mathscr{L}h^{1/2}d^{1/2})$
$\quad+h^{1/2}K^{1/2}\frac{\mathscr{L}^{2}d^{1/2}}{N^{1/2}}$
In order to ensure $\|\bar{\mu}_{K}-\mu_{*}\|_{\textsf{TV}}\leq\epsilon$, it suffices to choose
$h=\Theta\left(\frac{\mathscr{C}_{\textsf{LSI}}\epsilon^{2}}{\mathscr{L}^{3}d}\right),\quad K=\widetilde{\Theta}\left(\frac{\mathscr{L}^{3}d}{\mathscr{C}_{\textsf{LSI}}^{2}\epsilon^{2}}\right),\quad N=\Theta\left(\frac{\mathscr{L}^{4}d}{\mathscr{C}_{\textsf{LSI}}\epsilon^{2}}\right).$ (63)